### v2

# Sparse Hanson-Wright Inequalities with Applications

Abstract
We derive Hanson-Wright inequalities for the quadratic form of a random vector with sparse independent components. Specifically, we consider cases where the components of the random vector are sparse $\alpha$-subexponential random variables with $\alpha>0$. Our proof relies on a novel combinatorial approach to estimate the moments of the random quadratic form. In addition, we obtain a new Bernstein-type inequality for the sum of independent sparse $\alpha$-subexponential random variables. We present two applications with the sparse Hanson-Wright inequality: (1) Local law and complete eigenvector delocalization for sparse $\alpha$-subexponential Hermitian random matrices; (2) Concentration of the Euclidean norm for the linear transformation of a sparse $\alpha$-subexponential random vector.
11footnotetext: Department of Mathematics, University of California Irvine, Irvine, CA 92697, USA, yiyunh4@uci.edu.22footnotetext: Department of Mathematics, Hong Kong University of Science and Technology, Clear Water Bay, Kowloon, Hong Kong, kewang@ust.hk. 44footnotetext: Department of Mathematics, University of Southern California, 3620 Vermont Avenue Los Angeles, CA 90089, USA, yizhezhu@usc.edu.
1 Introduction
The Hanson-Wright-type inequalities concern the concentration of the quadratic forms of random vectors. More specifically, consider a random vector $X=(X_{1},\cdots,X_{n})^{T}$ with independent components and a deterministic matrix $A=(a_{ij})_{1\leq i,j\leq n}$. The primary focus is on understanding the concentration of the quadratic form
$S_{n}:=\sum_{i,j}a_{ij}X_{i}X_{j}=X^{T}AX,$
around its expectation $\mathbb{E}(S_{n})=\sum_{i,j}a_{ij}\mathbb{E}(X_{i}X_{j})=\mathbb{E}(X^{T}AX)$.
Hanson and Wright [23, 45] consider the case where $X_{i}$’s are centered sub-gaussian random variables. A random variable $\xi$ is sub-gaussian if $\mathbb{P}(|X-\mathbb{E}(X)|>t)\leq 2\exp(-t^{2}/K^{2})$ for any $t>0$. Its sub-gaussian norm is defined as
$\|\xi\|_{\Psi_{2}}:=\inf\{t>0:\mathbb{E}(\exp(|X|/t)^{2})\leq 2\}<\infty.$
Rudelson and Vershynin [35] provide a modern proof of the Hanson-Wright inequality, improving the dependence on the matrix operator norm in the tail bound from [23, 45]. Specifically, they show that if $\max_{1\leq i\leq n}\|X_{i}\|_{\Psi_{2}}\leq K$. Then for any $t\geq 0$,
$\mathbb{P}\left(|S_{n}-\mathbb{E}(S_{n})|\geq t\right)\leq 2\exp\left(-c\min\left\{\frac{t^{2}}{K^{4}\|A\|_{F}^{2}},\frac{t}{K^{2}\|A\|}\right\}\right).$ (1.1)
This paper aims to derive Hanson-Wright inequalities for random variables with a sparse structure. We specifically investigate the case where each $X_{i}=x_{i}\xi_{i}$, with $x_{i}\sim\mathrm{Ber}(p_{i})$ and all $x_{i},\xi_{i}$ being independent. The $\xi_{i}$’s are centered $\alpha$-subexponential random variables, characterized by a parameter $\alpha>0$ such that for any $t>0$, $\mathbb{P}(|\xi_{i}|\geq t)\leq 2\exp(-t^{\alpha}/K_{1}^{\alpha})$. Equivalent formulations for $\alpha$-subexponential random variables exist using the Orlicz norm $\|\xi_{i}\|_{\Psi_{\alpha}}<\infty$ or the $L_{r}$ norms: $\|\xi_{i}\|_{L_{r}}\leq K_{2}r^{1/\alpha}$. Note that $K_{1}$, $K_{2}$, and $\|\xi_{i}\|_{\Psi_{\alpha}}$ can be chosen to differ only by a constant factor (depending on $\alpha$). The class of $\alpha$-subexponential random variables includes sub-gaussian ($\alpha=2$), sub-exponential ($\alpha=1$), and heavier-tailed distributions ($\alpha\in(0,1)$). We allow $\alpha=\infty$, which corresponds to bounded random variables satisfying $|\xi_{i}|\leq K$ almost surely (or equivalently, $\|\xi_{i}\|_{L_{r}}\leq K$ for all $r\geq 1$). Our focus is on the sparse regime where $p_{i}\ll 1$.
Throughout this paper, $C,C^{\prime},c$ denote absolute constants that may vary in different equations, and $C_{\alpha}$ is an absolute constant depending only on $\alpha$.
Notations: Let $A=(a_{ij})$ be an $n\times n$ real matrix. The Frobenius norm of $A$ is defined as $\|A\|_{F}=\sqrt{\sum_{i,j}a_{ij}^{2}}$, and its operator norm is denoted by $\|A\|$. The maximum norm is given by $\|A\|_{\max}=\max_{i,j}|a_{ij}|$. We also denote $\|A\|_{1,\infty}=\max_{i}\sum_{j}|a_{ij}|$ and $\|A\|_{2,\infty}=\max_{i}(\sum_{j}|a_{ij}|^{2})^{1/2}$. For a vector $a\in\mathbb{R}^{n}$, $\|a\|$ denotes its Euclidean norm and $\|a\|_{\infty}$ denotes the maximum absolute value of its entries.
For a real random variable $X$, its $L_{r}$ ($r\geq 1$) norm is denoted by $\|X\|_{L_{r}}=(\mathbb{E}|X|^{r})^{1/r}$. The $\alpha$-Orlicz norm [^1] This is a quasi-norm for $\alpha\in(0,1).$ of a random variable $X$ is defined by
$\|X\|_{\Psi_{\alpha}}:=\inf\{t>0:\mathbb{E}(\exp(|X|/t)^{\alpha})\leq 2\}<\infty.$
For two functions $f(n),g(n)>0$, we use the asymptotic notations $g(n)\ll f(n)$ and $g(n)=o(f(n))$ if $f(n)/g(n)\to\infty$ as $n\to\infty$. Additionally, we use $f(n)\lesssim g(n)$ to denote that $f(n)\leq Cg(n)$ for some $C>0$.
To derive the concentration inequalities, we decompose the quadratic form $S_{n}$ into two parts: the diagonal part $S_{\mathrm{diag}}=\sum_{i=1}^{n}a_{ii}X_{i}^{2}$ and the off-diagonal part $S_{\mathrm{off}}=\sum_{i\neq j}a_{ij}X_{i}X_{j}$. We first establish concentration inequalities for $S_{\mathrm{off}}$, which presents the main technical challenge due to the dependencies among the terms in the sum. We then derive new concentration results for $S_{\mathrm{diag}}$, extending the existing theory for sums of independent random variables to the sparse setting.
THEOREM 1.
Let $X=(X_{1},\cdots,X_{n})^{T}$ be a random vector with independent centered sparse sub-exponential entries, i.e., $X_{i}=x_{i}\xi_{i}$ where $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\xi_{i}$ is a centered random variable satisfying $\|\xi_{i}\|_{L_{r}}\leq Kr$ for all $r\geq 1$ with some $K>0$. Let $A$ be an $n\times n$ diagonal-free matrix. Then there exists an absolute constant $C>0$ such that for any $t>0$,
$\mathbb{P}(|X^{T}AX|\geq t)\leq C^{\prime}\exp\left(-C\min\left\{\frac{t^{2}}{K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{K^{2}\gamma_{1,\infty}},\left(\frac{t}{K^{2}\|A\|_{\max}}\right)^{1/2}\right\}\right),$ (1.2)
where
$\gamma_{1,\infty}:=\max_{i}\left\{\sum_{j\neq i}|a_{ij}|p_{j},\sum_{j\neq i}|a_{ji}|p_{j}\right\}.$ (1.3)
REMARK 1.1.
We now discuss some generalizations of our results. The assumption that each $X_{i}$ is a product of independent Bernoulli and (centered) sub-exponential random variables can be relaxed in multiple ways. First, the conclusion of Theorem 1 holds for any independent centered random variables $X_{i}$ whose moments satisfy $\mathbb{E}|X_{i}|^{r}\leq p_{i}(Kr)^{r}$ for $r\geq 2$ and some $K>0$, where $p_{i}$ represents the sparsity parameter. Second, with a minor modification of our proof, the theorem extends to independent centered Bernoulli random variables, i.e., $X_{i}=x_{i}-p_{i}$ where $x_{i}\sim\mathrm{Ber}(p_{i})$. In this case, (1.2) holds with $K=1$. Finally, Theorem 1 generalizes to independent noncentered random variables $Y_{i}$, providing concentration results for $\sum_{i\neq j}a_{ij}Y_{i}Y_{j}$ around its mean $\sum_{i\neq j}a_{ij}\mathbb{E}(Y_{i})\mathbb{E}(Y_{j})$. This extension is detailed in Theorem 13 of Appendix B.
The tail bound in our theorem consists of three terms. The sub-gaussian tail is expected from the central limit theorem since
$\mathrm{Var}(X^{T}AX)=\sum_{i\neq j}(a_{ij}^{2}+a_{ij}a_{ji})p_{i}p_{j}\mathbb{E}(\xi_{i}^{2})\mathbb{E}(\xi_{j}^{2})\leq 32K^{4}\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}.$
The inequality above follows from $a_{ij}a_{ji}\leq(a_{ij}^{2}+a_{ji}^{2})/2.$ The sub-exponential tail dominates when $A$ has a prominent row or column. For instance, consider a case where only the first row of $A$ is non-zero, with non-negative entries for simplicity, and $X_{i}=x_{i}-p$ for all $i$ where $x_{i}\sim\mathrm{Ber}(p)$. Then $X^{T}AX=X_{1}\cdot(\sum_{i=2}^{n}a_{1i}X_{i})$ is the product of two independent random variables. If $p$ is chosen such that $p(\sum_{i=2}^{n}a_{1i})=p\|A\|_{1,\infty}=O(1)$, then $\sum_{i=2}^{n}a_{1i}x_{i}$ is approximately Poisson with parameter $p\|A\|_{1,\infty}$. The tail behavior of $X^{T}AX$ is controlled by $\exp(-ct/p\|A\|_{1,\infty})$. The $1/2$-subexponential tail emerges due to individual terms in the summation $X^{T}AX=\sum_{i\neq j}a_{ij}X_{i}X_{j}$. Conditioning on $x_{i}$’s, each $a_{ij}X_{i}X_{j}$ is a product of two independent sub-exponential random variables and has a tail of magnitude $\exp(-(t/K^{2}|a_{ij}|)^{1/2})$.
Moreover, for sparse $\alpha$-subexponential random vectors, we have the following generalization of Theorem 1.
THEOREM 2.
Consider a random vector $X=(X_{1},X_{2},\dots,X_{n})^{T}$ with $X_{i}=x_{i}\xi_{i}$ where all $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\xi_{i}$ is a centered random variable with $\|\xi_{i}\|_{L_{r}}\leq Kr^{\frac{1}{\alpha}}$, for any $r\geq 1$ and some $\alpha>0$. Assume $A$ is a diagonal-free $n\times n$ matrix. Then there exists a constant $C_{\alpha}>0$ depending only on $\alpha$ such that for any $t>0$,
$\mathbb{P}(|X^{T}AX|\geq t)\leq C^{\prime}\exp\left(-C_{\alpha}\min\left\{\frac{t^{2}}{K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{K^{2}\gamma_{1,\infty}},\left(\frac{t}{K^{2}\|A\|_{\max}}\right)^{\min\{\frac{\alpha}{2},\frac{1}{2}\}}\right\}\right).$ (1.4)
REMARK 1.2.
We see the third term in the tail bound of Theorem 2 does not gain any improvement for $\alpha>1$. It is an open question to see if Theorem 2 can be improved for $\xi_{i}$ being sub-gaussian. For sparse sub-gaussian random vectors, compared to (1.4), the inequality in [47] (presented later as Theorem 7) has $t/(K^{2}\|A\|)$ in the second term, which is independent of the sparsity $p_{i}$, but it does not involve the $t^{1/2}$ term.
For a general matrix $A$, to control the diagonal sum $S_{\mathrm{diag}}$, we introduce the following bound for the concentration of the sum of independent sparse $\alpha$-subexponential random variables:
THEOREM 3.
Consider a random vector $X=(X_{1},X_{2},\dots,X_{n})$ with $X_{i}=x_{i}\xi_{i}$ where all $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\xi_{i}$ is a centered random variable with $\|\xi_{i}\|_{r}\leq Kr^{\frac{1}{\alpha}}$, for all $r\geq 1$ and some $\alpha>0$. Let $a=(a_{1},\ldots,a_{n})$ be a deterministic vector in $\mathbb{R}^{n}$. Then for any $t>0$,
$\mathbb{P}\left(\left|\sum_{i=1}^{n}a_{i}X_{i}\right|\geq t\right)\leq C^{\prime}\exp\left(-C_{\alpha}\min\left\{\frac{t^{2}}{K^{2}\sum_{i}a_{i}^{2}p_{i}},\left(\frac{t}{K\|a\|_{\infty}}\right)^{\min\{\alpha,1\}}\right\}\right).$ (1.5)
Theorem 3 can be seen as a generalization of Bernstein’s inequality for $\alpha$-subexponential random variables with a dependence on the variance of the linear form. We present a brief discussion of related literature. Directly applying existing results from [22, Corollary 6.1], one will only get $K^{2}\sum_{i}a_{i}^{2}$ in the first denominator of (1.5) without the dependence on $p_{i}$. For sums of independent $\alpha$-subexponential random variables (termed sub-Weibull of order $\alpha$ in their work), [28, Theorem 3.1] established concentration bounds that, while not reflecting the sparsity $p_{i}$ in the sub-gaussian tail, achieve a sharper second term $\exp(-(t/KL_{\alpha}(a))^{\alpha})$, where $L_{\alpha}(a)=\|a\|_{\infty}$ if $\alpha\leq 1$ and $L_{\alpha}(a)=\|a\|_{\frac{\alpha}{\alpha-1}}$ if $\alpha>1$. Regarding variance-scaled tail bounds, [28, Theorem 3.2-3.3] proved the following tail bound:
$2\exp\left(-\min\left\{C\frac{t^{2}}{\sum_{i}a_{i}^{2}p_{i}\mathbb{E}\xi_{i}^{2}},C_{\alpha}\left(\frac{t}{K\|a\|_{\infty}\log^{\frac{1}{\alpha}}(n+1)}\right)^{\min\{\alpha,1\}}\right\}\right).$
The factor $(\log(n+1))^{1/\alpha}$ is shown to be sharp by [1, Section 2.2], assuming only $\|X_{i}\|_{\psi_{\alpha}}<\infty$ and variance-based norm bounds. In contrast, our new bound (1.5) captures the sparsity parameters $p_{i}$ in the sub-gaussian tail with a variance proxy $K^{2}\sum_{i}a_{i}^{2}p_{i}$, while maintaining a dimension-free second term.
We now bound the concentration of the diagonal sum $S_{\mathrm{diag}}$ in the quadratic form $\sum_{i,j}a_{ij}X_{i}X_{j}$ by applying Theorem 3 to $X_{i}^{2}-EX_{i}^{2}$, which are sparse $\frac{\alpha}{2}$-subexponential random variables.
COROLLARY 1.
Consider a random vector $X=(X_{1},X_{2},\dots,X_{n})$ with $X_{i}=x_{i}\xi_{i}$ where all $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\xi_{i}$ is a centered random variable with $\|\xi_{i}\|_{r}\leq Kr^{\frac{1}{\alpha}}$, for all $r\geq 1$ and some $\alpha>0$. Then for any $t>0$,
$\mathbb{P}\left(\left|\sum_{i=1}^{n}a_{ii}(X_{i}^{2}-\mathbb{E}X_{i}^{2})\right|\geq t\right)$ (1.6)
$\leq$ $\,C^{\prime}\exp\left(-C_{\alpha}\min\left\{\frac{t^{2}}{K^{4}\sum_{i}a_{ii}^{2}p_{i}},\left(\frac{t}{K^{2}\max_{i}|a_{ii}|}\right)^{\min\{\frac{\alpha}{2},1\}}\right\}\right).$ (1.7)
Combining the bounds for the diagonal and off-diagonal terms, we have the following result for random quadratic forms of a general matrix $A$:
COROLLARY 2.
Let $X=(X_{1},\cdots,X_{n})^{T}$ be a random vector with independent centered sparse $\alpha$-subexponential entries, i.e., $X_{i}=x_{i}\xi_{i}$ where $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\xi_{i}$ is a centered random variables with $\|\xi_{i}\|_{L_{r}}\leq Kr^{\frac{1}{\alpha}}$ for any $r\geq 1$ and some $\alpha>0$. Let $A$ be an $n\times n$ matrix. Then for any $t>0$,
$\mathbb{P}\left(\left|X^{T}AX-\mathbb{E}(X^{T}AX)\right|\geq t\right)$
$\leq$ $C^{\prime}\exp\left(-C_{\alpha}\min\left\{\frac{t^{2}}{K^{4}(\sum_{i}a_{ii}^{2}p_{i}+\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j})},\frac{t}{K^{2}\gamma_{1,\infty}},\left(\frac{t}{K^{2}\|A\|_{\max}}\right)^{\min\{\frac{\alpha}{2},\frac{1}{2}\}}\right\}\right),$
where
$\gamma_{1,\infty}:=\max_{i}\left\{\sum_{j\neq i}|a_{ij}|p_{j},\sum_{j\neq i}|a_{ji}|p_{j},|a_{ii}|\right\}.$ (1.8)
Note that the definition of $\gamma_{1,\infty}$ in (1.8) is consistent with the diagonal-free case in (1.3). Furthermore, Theorem 2, Corollary 1, and Corollary 2 can be extended through generalizations similar to that of Theorem 1, as discussed in Remark 1.1. We omit the details for brevity.
The proof of Hanson-Wright inequalities for sub-gaussian random vectors in [35] replies on bounding the exponential generating function with a Gaussian comparison argument. Due to the sparsity parameters, a direct adaptation of the proof in [35] will yield a dependence on the sub-gaussian norm of $x_{i}\xi_{i}$, which does not have the optimal dependence on $p_{i}$. Moreover, this technique does not apply to $\alpha$-subexponential random variables. Instead, our proof is based on the moment method, which captures the correct variance term in the sub-gaussian tail part and the sub-exponential tail part also depends on the sparsity parameters. The moment method approach was used in [24, 37, 38]. Our inequality improves the one in [24] to have sub-gaussian and sub-exponential tails, and generalize the inequalities from [37] to sparse $\alpha$-subexponential random vectors with a simpler proof.
We present two applications of the new sparse Hanson-Wright inequality. The first application is a new local law and eigenvector delocalization for Hermitian random matrices with independent sparse $\alpha$-subexponential entries. The second is the concentration of the norm of the linear transformation of an $\alpha$-sparse subexponential random vector.
Organization: The rest of the paper is organized as follows. Section 2 reviews existing literature on sparse Hanson-Wright inequalities, along with a discussion and comparison of our findings with previous work. We present two applications of our sparse Hanson-Wright inequality in Section 3. The proofs of our main theoretical results are provided in Sections 4 and 5, while Section 6 contains the proofs of the applications. Additional technical results and generalizations of our theorems are presented in Appendices A and B, respectively. Finally, Appendices C and D provide the proofs of auxiliary results used in the applications.
2 Previous results and comparison
Recent progress in the Hanson-Wright-type inequalities and more generally, the polynomials of independent random variables, can be found in [42, 2, 47, 24, 46, 27, 48, 31, 15, 22, 36, 6, 26, 38, 16, 36, 22, 13, 43] and references therein. Specifically, an extension to $\alpha$-subexponential random vectors was obtained in [22]:
THEOREM 4 (Corollary 1.4 (2) in [22]).
For any $\alpha\in(0,1]$, let $X_{1},\dots,X_{n}$ be independent centered random variables with $\|X_{i}\|_{\Psi_{\alpha}}\leq K$, and $A$ be a symmetric matrix. Then for any $t\geq 0$,
$\mathbb{P}$ $\left(|X^{T}AX-\mathbb{E}X^{T}AX|\geq t\right)$
$\leq 2\exp\left(-{C_{\alpha}}\min\left\{\frac{t^{2}}{K^{4}\|A\|_{F}^{2}},\frac{t}{K^{2}\|A\|},\left(\frac{t}{K^{2}\|A\|_{2,\infty}}\right)^{\frac{2\alpha}{2+\alpha}},\left(\frac{t}{K^{2}\|A\|_{\max}}\right)^{\alpha/2}\right\}\right).$ (2.1)
The above tail bound encompasses up to four different regimes. It represents a refined version of the analog to the Hanson-Wright inequality (for large $t$), as presented in Proposition 1.1 of [22]:
$\mathbb{P}$ $\left(|X^{T}AX-\mathbb{E}X^{T}AX|\geq t\right)\leq 2\exp\left(-{C_{\alpha}}\min\big{\{}{t^{2}}/{K^{4}\|A\|_{F}^{2}},\left({t}/{K^{2}\|A\|}\right)^{\alpha/2}\big{\}}\right).$ (2.2)
The extension of (2.2) to any $\alpha\in(0,2]$ is discussed in Sambale’s work [36].
The concentration of multivariate polynomials of independent random variables has been a subject of study since the work of Kim and Vu [26]. Schudy and Sviridenko [37] derived concentration inequalities for polynomials of independent central moment bounded random variables, improving upon their earlier findings in [38]. A random variable $Z$ is considered central moment bounded if $\mathbb{E}(|Z-\mathbb{E}Z|^{i})\leq iL\mathbb{E}(|Z-\mathbb{E}Z|^{i-1})$ for any integer $i\geq 1$. This class of central moment bounded random variables encompasses bounded and log-concave random variables, with Gaussian and exponential distributions as special examples.
The results by Schudy and Sviridenko [37], when applied specifically to quadratic forms of sparse exponential random variables, yield the following:
THEOREM 5 (Special case of Theorem 1.3 in [37]).
Let $X_{1},\cdots,X_{n}$ be independent centered sparse exponential random variables. Each $X_{i}=x_{i}\eta_{i}$ where $x_{i},\eta_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\eta_{i}$ is a centered $\mathrm{Exp}(K^{-1})$ random variable. Let $A$ be an $n\times n$ symmetric, diagonal-free matrix. For any $t>0$,
$\mathbb{P}\left(\left|X^{T}AX\right|\geq t\right)\leq 2\exp\left(-C\min\left\{\frac{t^{2}}{K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{K\gamma_{1,\infty}},\left(\frac{t}{K^{2}\|A\|_{\max}}\right)^{1/2}\right\}\right).$ (2.3)
This result aligns with our Theorem 1 for the sparse exponential random variables. However, our findings extend beyond this, encompassing sparse sub-exponential random variables and further generalizing to sparse $\alpha$-subexponential random variables, as given in Theorem 2. Our proof is also simpler compared to [38]. It is worth noting that, similar to the discussion in [22, Section 1.2], a central moment bounded random variable $Z$ can be shown to be sub-exponential, i.e., $\|Z\|_{\Psi_{1}}<\infty$. The converse implication is not immediately evident and may not hold in all cases.
Another line of research focuses on the tail bounds of (unbounded) U-statistics. Chakrabortty and Kuchibhotla [14] examine a particular class of U-statistics. These U-statistics take the form: $U_{n}=\sum_{1\leq i\neq j\leq n}\phi(Y_{i})w_{ij}(X_{i},X_{j})\phi(Y_{j})$, where $Z_{1}=(X_{1},Y_{1}),\cdots,Z_{n}=(X_{n},Y_{n})$ are independent random vectors. When applied to sparse $\alpha$-subexponential random variables, their results yield the following moment estimates:
THEOREM 6 (Theorem 1 from [14]).
Under the same assumptions as Theorem 2, denote $\alpha^{*}:=\min\{\alpha,1\}$. For any integer $r\geq 1$,
$\|X^{T}AX\|_{L^{r}}\lesssim$ $\sqrt{r}\cdot K_{\alpha}^{2}p\|A\|_{F}+r\cdot K_{\alpha}^{2}p\|A\|$
$+r^{\frac{1}{\alpha^{*}}}\cdot K_{\alpha}^{2}\left[(\log n)^{\frac{1}{2}+\frac{1}{\alpha}}\sqrt{p}\|A\|_{2,\infty}+(\log n)^{1+\frac{2}{\alpha}}\|A\|_{\max}\right]$
$+r^{\frac{1}{2}+\frac{1}{\alpha^{*}}}\cdot K_{\alpha}^{2}(\log n)^{\frac{1}{\alpha}}\sqrt{p}\|A\|_{2,\infty}+r^{\frac{2}{\alpha^{*}}}\cdot K_{\alpha}^{2}(\log n)^{\frac{2}{\alpha}}\|A\|_{\max}.$ (2.4)
The bound above comprises a mixture of four terms. Notably, for $\alpha\in(0,1]$, the tail bound implied by these moment estimates is consistent with the result in Theorem 4 for constant $p$, up to the $\log n$ factors.
An intriguing future direction would be to derive a refined concentration inequality that encompasses both our result (Theorem 2) and the above results. Recent work by [7] has extended the results in [14], generalizing their findings to a broader class of unbounded, heavy-tailed U-statistics.
There is also considerable interest in the study of quadratic forms involving sparse sub-gaussian random variables [47, 6, 34]. A key result was obtained by Zhou [47]:
THEOREM 7 (Theorem 1.1 from [47]).
Let $X_{1},\cdots,X_{n}$ be independent centered sparse sub-gaussian random variables, i.e. $X_{i}=x_{i}\xi_{i}$ where $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\xi_{i}$ is a centered sub-gaussian random variables with $\|\xi_{i}\|_{\Psi_{2}}\leq K$. Let $A$ be an $n\times n$ matrix. Denote $S_{n}=\sum_{i,j}a_{ij}x_{i}x_{j}\xi_{i}\xi_{j}$. Then for any $t>0$,
$\mathbb{P}\left(\left|S_{n}-\mathbb{E}(S_{n})\right|>t\right)\leq 2\exp\left(-c\min\left\{\frac{t^{2}}{K^{4}(\sum_{k=1}^{n}p_{k}a_{kk}^{2}+\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j})},\frac{t}{K^{2}\|A\|}\right\}\right).$
Our Corollary 2, when applied to $\alpha=2$, reveals up to three different regimes, in contrast to the two regimes as above. A comparable bound to [47] is derived in [34] for the generalized setting $S=\sum_{i,j}a_{ij}x_{1i}x_{2j}\xi_{1i}\xi_{2j}$ where the pairs $(x_{1i},x_{2i})$ and $(\xi_{1i},\xi_{2i})$ are independent, although $x_{1i},x_{2i}$ or $\xi_{1i},\xi_{2i}$ may be correlated. Furthermore, [6] has recently established a result similar to [47], under the assumptions that $\|A\|\leq 1$ and all Bernoulli random variables share the same parameter $p$. The concentration inequality in Lemma 2.8 from [6] yields a sharp bound assuming $1\gg p\gg\log n/n$.
The concentration of quadratic forms involving (sparse) Bernoulli random variables, specifically $\sum_{i,j}a_{ij}x_{i}x_{j}$ where $x_{i}$’s are independent Bernoulli (or centered Bernoulli) random variables, has been a subject of extensive study in the literature. While not the primary focus of our current work, we find it valuable to offer a concise summary of selected results in this area.
For simplicity, for the rest of this section, we assume that $A$ is symmetric, diagonal-free, and $x_{i}$ are independent centered Bernoulli random variables with parameter $p_{i}$. Hence, $\sum_{i\neq j}a_{ij}x_{i}x_{j}:=x^{T}Ax$ is centered.
Latała [29] provides precise moment estimates for the quadratic forms of Rademacher random variables, where all $p_{i}=\frac{1}{2}$. For any $r\geq 1$,
$C^{-1}(\|A\|_{\{1,2\},r}+\|A\|_{\{1\},\{2\},r})\leq\|x^{T}Ax\|_{L^{r}}\leq C(\|A\|_{\{1,2\},r}+\|A\|_{\{1\},\{2\},r}).$ (2.5)
The definitions of matrix norms above are involved and can be found in [29] or [3, Section 5.2]. These norms satisfy that $\|A\|_{\{1,2\},r}\leq\sqrt{r}\|A\|_{F}$ and $\|A\|_{\{1\},\{2\},r}\leq r\|A\|$, which allows this result to recover the Hanson-Wright inequality (1.1).
Giné, Latała, and Zine’s study [21] on the concentration of bounded U-statistics offers insights that are useful for deriving tail estimates of random quadratic forms. By applying decoupling results and performing straightforward calculations, one can obtain the following tail estimate as a special case of their Theorem 3.3 in [21]: for any $t>0$,
$\mathbb{P}\left(\left|x^{T}Ax\right|\geq t\right)\leq 2\exp\left(-C\min\left\{\frac{t^{2}}{\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{B},\left(\frac{t}{D}\right)^{2/3},\left(\frac{t}{\|A\|_{\max}}\right)^{1/2}\right\}\right),$ (2.6)
where $B=(\max_{i}p_{i})\|A\|$ and $D=\max_{j}\sqrt{\sum_{i}a_{ij}^{2}p_{i}}$. The $t^{2}$ and $t^{1/2}$ terms in (2.6) match our Theorem 1 for bounded random variables. The $t$ term in some cases (e.g., all $p_{i}$ are the same) can be better than Theorem 1, while Theorem 1 does not have the $t^{2/3}$ term.
The result by Schudy and Sviridenko [37], as mentioned earlier, applies to the sparse Bernoulli random variable case. Their findings yield the tail bound:
$\mathbb{P}(|x^{T}Ax|\geq t)\leq 2\exp\left(-C\min\left\{\frac{t^{2}}{\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{\max_{i}\sum_{j}|a_{ij}|p_{j}},\left(\frac{t}{\|A\|_{\max}}\right)^{1/2}\right\}\right).$ (2.7)
Note that a direct adaptation of our proof for Theorem 1 provides the same bound under this setting.
Bellec [8] derived a sharper version of the Hanson-Wright inequality than (1.1) for random variables satisfying a Bernstein moment condition. This improved inequality captures the variance of $X^{T}AX$ in the sub-gaussian tail. Their result applies to sparse Bernoulli random variables. He, Knowles, and Marcozzi [24] established an upper bound on the $L_{r}$ norms $\|x^{T}Ax\|_{L^{r}}$, assuming all $p_{i}=p$. Notably, their bound involves $\|A\|_{2,\infty}$ and $\|A\|_{\max}$. This estimate is crucial in analyzing the spectral statistics of sparse Erdős-Rényi random graphs $G(n,p)$ at the critical scale $pn=O(\log n)$. Zhou [48] (see Theorem 2.5) examined the quadratic forms of non-identical independent Bernoulli random variables, providing bounds on the moment generating function (MGF) of the off-diagonal term.
Addressing the asymptotic behavior, [12] established the limiting distributions for the random quadratic form $S_{n}=\sum_{i,j}a_{ij}x_{i}x_{j}$, where $A=(a_{ij})$ is a symmetric, $\{0,1\}$-valued, diagonal-free matrix and $x_{i}$ are i.i.d. Bernoulli random variables with parameter $p$. Their study focused on the sparse regime where $\mathbb{E}(S_{n})=p^{2}\sum_{i,j}a_{ij}=O(1)$.
We conclude the literature review with two open questions. It is unclear to us if a sparsity-adaptive Hanson-Wright inequality exists for Bernoulli random vectors that is optimal for all sparsity regimes. Taking $p_{i}$ to be a constant independent of $n$ in (2.6) and (2.7) does not recover the Hanson-Wright inequality for sub-gaussian random vectors in [35]. And (2.6) and (2.7) yields different bounds depending on various norms of $A$. For sparse sub-gaussian random vectors, it would be interesting to unify our Theorem 1 and Zhou’s inequality in [47].
3 Applications
3.1 Sparse Wigner matrices: local law and eigenvector delocalization
In this section, we apply concentration inequalities developed in this paper to study the spectral statistics of sparse Wigner matrices with $\alpha$-subexponential entries for $\alpha>0$. Specifically, we consider a symmetric matrix $H\in\mathbb{R}^{n\times n}$ with sparse $\alpha$-subexponential entries defined as
$H=\frac{1}{\sqrt{np}}\widetilde{H}=\frac{1}{\sqrt{np}}(\widetilde{H}_{ij}),$ (3.1)
where $\widetilde{H}_{ij}$ ($1\leq i\leq j\leq n$) are independent centered random variables. Each entry $\widetilde{H}_{ij}$ has the form $\widetilde{H}_{ij}=x_{ij}\xi_{ij}$, where $x_{ij}$ and $\xi_{ij}$ are independent, with $x_{ij}\sim\mathrm{Ber}(p)$ and $\xi_{ij}$ having mean 0, variance 1, and being $\alpha$-subexponential. We assume that for any $r\geq 1$, $\max_{i,j}\|\xi_{ij}\|_{L_{r}}\leq Kr^{\frac{1}{\alpha}}$, where $K\geq 1$ is a constant. Here $p=p(n)\in(0,1)$.
Denote the eigenvalues of $H$ as $\lambda_{1}(H),\cdots,\lambda_{n}(H)$ and their corresponding normalized eigenvectors as $u_{1},\cdots,u_{n}$. When $np\to\infty$, it is known that the empirical spectral distribution (ESD) of $H$, defined as $\mu_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{\lambda_{i}(H)}$, converges weakly to the semicirle law in probability [44, 41]. When $\xi_{ij}=1$, $H$ is the adjacency matrix of Erdős-Rényi graph, and its spectral properties, including the local law and eigenvector delocalization [41, 24], universality [18, 19, 25], and extreme eigenvalues [9, 5], have received considerable attention in recent years. When $\xi_{ij}$ is a bounded random variable, several extensions have been obtained in [5, 40, 6, 17]. When $p=\frac{c}{n}$, large deviation principle of the spectral edge was studied in [20].
We establish the local law and complete eigenvector delocalization for $H$, analogous to He, Knowles, and Marcozzi’s work [24] that studied the (normalized) adjacency matrix of the Erdős-Rényi random graph $G(n,p)$ in the supercritical regime $np\geq C\log n$. Our methodology follows the framework of [24], a method that was developed through a series of works [19, 18] (see references therein) and is comprehensively surveyed in [10]. The key innovation in our approach lies in the application of the large deviation inequalities developed in this paper to handle the sparse $\alpha$-subexponential setting.
The main tool to study the spectral statistics of $H$ is via its Green function
$G(z):=(H-z)^{-1}$
for $z\in\mathbb{C}\setminus\{\lambda_{1}(H),\ldots,\lambda_{n}(H)\}$. Here we abbreviate $(H-zI_{n})^{-1}$ as $(H-z)^{-1}$. Define the spectral domain
$\mathbf{S}:=\{E+\mathrm{i}\eta\in\mathbb{C}:|E|\leq 10,n^{-1}<\eta\leq 1\}.$ (3.2)
For $\operatorname{Im}z\neq 0$, the Stieltjes transform $m(z)$ of the semicircle law is defined as
$m(z):=\int\frac{\rho_{sc}(dx)}{x-z},\quad\rho_{sc}(dx):=\frac{1}{2\pi}\sqrt{(4-x^{2})_{+}}dx.$
Furthermore, $m(z)$ is characterized as the unique solution of the quadratic equation
$m(z)+\frac{1}{m(z)}+z=0$ (3.3)
such that $\operatorname{Im}m(z)>0$ for $\operatorname{Im}z>0$. We derive the following local law that estimates the matrix elements of $G(z)$ for $z\in\mathbf{S}$. Assume $n\geq 10$.
THEOREM 8 (Local law for sparse Wigner matrix with $\alpha$-subexponential entries).
Consider the matrix $H$ in (3.1) and its Green function $G(z)$. For any $D>0$ and $\delta\in(0,1)$, there exists $L\equiv L(\delta,D)$ such that the following holds. Let $p\geq LC_{\alpha}^{2}K^{4}\frac{(\log n)^{\max\{2,\frac{2}{\alpha}\}}}{n}$. For $z=E+\mathrm{i}\eta\in\mathbf{S}$ with $\eta\geq LC_{\alpha}^{2}K^{4}\frac{\log^{2}n}{n}$, we have
$\mathbb{P}\left(\max_{i,j}\left|G_{ij}(z)-m(z)\delta_{ij}\right|\leq\delta\right)\geq 1-n^{-D}.$ (3.4)
Here, $C_{\alpha}\geq 1$ is an absolute constant depending only on $\alpha$.
A standard consequence of the local law is the semicircle law on small scales given below. Recall that $\mu_{n}$ is the ESD of $H$.
THEOREM 9 (Semicircle law on small scales).
Under the assumption of Theorem 8, for any $D>0$ and interval $I\subset[-5,5]$ satisfying $|I|\geq LC_{\alpha}^{2}K^{4}\frac{\log^{2}n}{n}$, we have
$\mathbb{P}\left(|\mu_{n}(I)-\rho_{sc}(I)|\leq\delta|I|\right)\geq 1-n^{-D}.$
The proof of Theorem 9 can be derived by either following the proof detailed in Section 8 from [10] or applying Lemma 64 from [39] together with Theorem 8. We omit the details.
Another consequence of the local law is the complete delocalization of the (normalized) eigenvectors $\{u_{i}\}_{i=1}^{n}$ of $H$.
THEOREM 10 (Complete eigenvector delocalization).
Consider the matrix $H$ in (3.1). For any $D>0$, there exist $L=L(D)$ and $n_{0}=n_{0}(D)$ such that for $p\geq LC_{\alpha}^{2}K^{4}\frac{(\log n)^{1+\max\{\frac{2}{\alpha},1\}}}{n}$ and $n\geq n_{0}$, we have
$\mathbb{P}\left(\forall i:\|u_{i}\|_{\infty}\leq\sqrt{2L}C_{\alpha}K^{2}\frac{\log n}{\sqrt{n}}\right)\geq 1-2n^{-D}.$
Here, $C_{\alpha}\geq 1$ is an absolute constant depending only on $\alpha$.
The study of the complete eigenvector delocalization has a rich history (see survey [33]). For the adjacency matrix of the Erdős-Rényi random graph $G(n,p)$, [24] showed that all eigenvectors satisfy $\|u_{i}\|_{\infty}\leq\frac{n^{(\log n)^{-1/2}}}{\sqrt{n}}$ with high probability when $p\geq C\frac{\sqrt{\log n}}{\sqrt{n}}$, improving upon the bounds established in [41]. The conjectured optimal bound is $\|u_{i}\|_{\infty}\lesssim\frac{\sqrt{\log n}}{\sqrt{n}}$, which has been proven for bulk eigenvectors of Wigner matrices with bounded entries [42]. For all eigenvectors, the optimal delocalization for Wigner matrices with sub-exponential entries was established in [11]. Our bound in Theorem 10 appears to be $\sqrt{\log n}$ away from this optimal bound in the dense case, yet represents the current best result for Wigner matrices with sparse $\alpha$-subexponential entries. The proof of Theorem 10 follows by combining two key ingredients: the local law and an estimate of the spectral norm of $H$. The complete details are presented in Appendix C.
By the same proof method with the sparse Hanson-Wright inequality for Bernoulli random vectors in (2.3), we can obtain $\|u_{i}\|_{\infty}\lesssim\frac{\log n}{\sqrt{n}}$ for $p\geq\frac{C\log^{2}n}{\sqrt{n}}$, which improves the delocalization estimate of [24] in the super-critical regime $p\geq\frac{C\log^{2}n}{\sqrt{n}}$.
REMARK 3.1.
The results presented in this section can be extended in several directions. First, the matrix $H$ can be complex-valued, requiring only cosmetic modifications to our proofs. The entries may satisfy more general moment conditions as discussed in Remark 1.1. Moreover, the diagonal entries can have a common variance that differs from the variance of the off-diagonal entries. A more substantial generalization involves sparse Wigner-type matrices with heterogeneous sparsity parameters $p_{ij}$, as studied in [17]. Such extensions can be achieved using the quadratic vector equation developed in [4]. While these generalizations are natural and tractable, we do not pursue their details in this work.
3.2 Concentration of the Euclidean norm of sparse random vectors
Another application of our sparse Hanson-Wright inequality concerns the estimation of the Euclidean norm of random vectors, an application previously considered in [35, 22] in the non-sparse setting. For a sparse $\alpha$-subexponential random vector $X$ and a deterministic matrix $B$, we establish the following concentration result for the squared norm $\|BX\|^{2}$.
THEOREM 11.
Consider $X=(X_{1},\dots,X_{n})^{T}$ with $X_{i}=x_{i}\xi_{i}$ where all $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p)$ and $\xi_{i}$ is a mean-zero, unit-variance random variable with $\|\xi_{i}\|_{L_{r}}\leq Kr^{\frac{1}{\alpha}}$ for all $r\geq 1$ and some $\alpha>0$. Let $B\in\mathbb{R}^{m\times n}$ be an deterministic matrix with columns $B_{1},\dots,B_{n}$. For any $s>0$, with probability at least $1-C^{\prime}\exp(-cs)$, the following holds:
$\left|\|BX\|^{2}-p\|B\|_{F}^{2}\right|\leq$ $\sqrt{s}K^{2}\sqrt{p^{2}\|B^{T}B\|_{F}^{2}+p\sum_{i}\|B_{i}\|^{2}}$ (3.5)
$+sK^{2}\max_{1\leq i\leq n}\Big{\{}p\sum_{j\neq i}|\langle B_{i},B_{j}\rangle|,\|B_{i}\|^{2}\Big{\}}+s^{\max\{\frac{2}{\alpha},2\}}K^{2}\|B^{T}\|_{2,\infty}^{2}.$ (3.6)
Note that in the above result, $\|B^{T}\|_{2,\infty}=\max_{i}\|B_{i}\|$ denotes the maximum Euclidean norm among all columns of $B$.
Let us compare our result with the existing bound in [22] for the special case $\alpha=1$, although both their result and ours apply to $\alpha\in(0,1]$. Consider a sparse sub-exponential random vector $X$ with i.i.d. entries $X_{i}=x_{i}\xi_{i}$ where all $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p)$ and $\xi_{i}$ is a mean-zero, unit-variance random variable with $\|\xi_{i}\|_{L_{r}}\leq Kr$ for all $r\geq 1$. These conditions imply $\|X_{i}\|_{\Psi_{1}}\leq CK$ for a universal constant $C>0$. According to [22, Corollary 2.2], with probability at least $1-C^{\prime}\exp(-cs)$,
$\left|\|BX\|^{2}-p\|B\|_{F}^{2}\right|\leq$ $\sqrt{s}K^{2}\|B^{T}B\|_{F}+sK^{2}\|B\|^{2}$ (3.7)
$+s^{3/2}K^{2}\max_{i}\sqrt{\sum_{j}\langle B_{i},B_{j}\rangle^{2}}+s^{2}K^{2}\|B^{T}\|_{2,\infty}^{2}.$ (3.8)
Compared to (3.8), our bound (3.6) does not have the $s^{3/2}$ term, while the $s^{2}$ terms in both bounds are identical. For small $p$, the $\sqrt{s}$ term in (3.6) improves upon the corresponding term in (3.8) since $\sqrt{\sum_{i}\|B_{i}\|^{2}}\leq\|B^{T}B\|_{F}$. When $p$ is sufficiently small, the $s$ term in (3.6) also improves the corresponding term in (3.8) due to the fact that $\max_{i}\|B_{i}\|\leq\|B\|$.
4 Proof of Theorem 1
The key input of the proof of Theorem 1 is the following moment estimates, whose proof is deferred to Section 4.1.
LEMMA 4.1.
Let $r$ be an even integer. Let $X_{1},\cdots,X_{n}$ be independent centered sparse sub-gaussian random variables, i.e. $X_{i}=x_{i}\xi_{i}$, where $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\|\xi_{i}\|_{L_{r}}\leq Kr$ for all $r\geq 1$ with a constant $K>0$. Let $A$ be a diagonal-free $n\times n$ matrix. Then
$\mathbb{E}\left|\sum_{i\neq j}a_{ij}X_{i}X_{j}\right|^{r}\leq(CK^{2})^{r}\max\left\{\left(r\sum_{i,j}a_{ij}^{2}p_{i}p_{j}\right)^{r/2},{(r\gamma_{1,\infty})^{r}},{(r^{2}\|A\|_{\max})^{r}}\right\}.$ (4.1)
With Lemma 4.1, we now proceed to prove our main result.
Proof of Theorem 1.
Combining the moment bound from Lemma 4.1 and Markov’s inequality, we derive
$\mathbb{P}\left(\left|\sum_{i\not=j}a_{ij}X_{i}X_{j}\right|\geq t\right)$ $\leq\frac{\mathbb{E}\left|\sum_{i\not=j}a_{ij}X_{i}X_{j}\right|^{r}}{t^{r}}$ (4.2)
$\leq\left(\frac{1}{t}\max\left\{CK^{2}\sqrt{r\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},CK^{2}r\gamma_{1,\infty},CK^{2}r^{2}\|A\|_{\max}\right\}\right)^{r}.$ (4.3)
For any $t>0$, if
$\min\left\{\frac{t^{2}}{Ce^{2}K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{CeK^{2}\gamma_{1,\infty}},\sqrt{\frac{t}{CeK^{2}\|A\|_{\max}}}\right\}<2,$ (4.4)
then in the inequality in Theorem 1 there is
$\exp\left(-\min\left\{\frac{t^{2}}{Ce^{2}K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{CeK^{2}\gamma_{1,\infty}},\sqrt{\frac{t}{CeK^{2}\|A\|_{\max}}}\right\}\right)\geq e^{-2}.$
We can choose constant $C^{\prime}=e^{2}$ such that the right-hand side in Theorem 1 is at least 1 and the theorem holds.
Next we assume the (4.4) fails. By choosing $r$ to be the largest positive even integer with
$2\leq r\leq\min\left\{\frac{t^{2}}{Ce^{2}K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{CeK^{2}\gamma_{1,\infty}},\sqrt{\frac{t}{CeK^{2}\|A\|_{\max}}}\right\},$
we have that
$\frac{1}{t}\max\left\{CK^{2}\sqrt{r\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},CK^{2}r\gamma_{1,\infty},CK^{2}r^{2}\|A\|_{\max}\right\}\leq\frac{1}{e},$
and further
$\mathbb{P}\left(\left|\sum_{i\not=j}a_{ij}X_{i}X_{j}\right|\geq t\right)$ $\leq e^{-r}$
$\leq\exp\left(-\left(c\min\left\{\frac{t^{2}}{K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{K^{2}\gamma_{1,\infty}},\sqrt{\frac{t}{K^{2}\|A\|_{\max}}}\right\}-2\right)\right)$
$=e^{2}\exp\left(-c\min\left\{\frac{t^{2}}{K^{4}\sum_{i,j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{K^{2}\gamma_{1,\infty}},\sqrt{\frac{t}{K^{2}\|A\|_{\max}}}\right\}\right).$
4.1 Proof of Lemma 4.1
In the remainder of this paper, we use $C$ to represent an absolute constant whose value may change from line to line. We do not keep track of these absolute constants explicitly.
Without loss of generality, we can assume $a_{ij}\geq 0$, as the right-hand side in Lemma 4.1 remains the same if we replace every $a_{ij}$ with $|a_{ij}|$. Moreover, note that
$X^{T}AX=X^{T}\frac{1}{2}(A+A^{T})X:=X^{T}\widetilde{A}X^{T},$
where $\widetilde{A}=(\widetilde{a}_{ij})$ is symmetric and diagonal-free. It is straightforward to verify that $\sum_{i,j}\widetilde{a}_{ij}^{2}p_{i}p_{j}\leq\sum_{i,j}a_{ij}^{2}p_{i}p_{j}$, $\max_{i}\sum_{j\neq i}|\widetilde{a}_{ij}|p_{j}\leq\gamma_{1,\infty}$, and $\|\widetilde{A}\|_{\max}\leq\|A\|_{\max}$. Therefore, it suffices to assume $A$ is symmetric in the subsequent proof.
Step 1. (Rewrite the moment). We start with the expansion
$\mathbb{E}\left|\sum_{i,j=1}^{n}a_{ij}X_{i}X_{j}\right|^{r}=\sum_{i_{1},\dots,i_{2r}}a_{i_{1}i_{2}}\cdots a_{i_{2r-1}i_{2r}}\mathbb{E}[X_{i_{1}}\cdots X_{i_{2r}}].$ (4.5)
Since $X_{i}$ are centered random variables, for the expectations to be nonzero, every index $i_{l}$ in the summation on the right-hand side of (4.5) must appear at least twice. We introduce the following notation.
  • Let $\mathcal{P}(2r)$ be the set of partitions of $[2r]$, and let $\mathcal{P}_{\geq 2}(2r)$ be the set of partitions of $[2r]$ whose blocks are of size at least $2$. For each $\Pi\in\mathcal{P}_{\geq 2}(2r)$, let $s\in[n]^{|\Pi|}$ assign indices $[n]$ to the blocks in $\Pi$. Denote $\mathcal{P}_{\geq 2}(2r,k)\subset\mathcal{P}_{\geq 2}(2r)$ to be the set of partitions of $[2r]$ with exactly $k$ blocks, each of size at least 2. Hence, we focus on $2\leq k\leq r$ since $A$ is diagonal-free and $k=1$ gives a zero in the sum.
  • Encode each $\Pi\in\mathcal{P}_{\geq 2}(2r,k)$ as a multigraph $G(\Pi)$ whose vertex set is $\Pi$. More precisely, consider $\mathcal{G}=([2r],E)$ to be the graph with vertex set $[2r]$ and edge set
    $\{\{1,2\},\{3,4\},\cdots,\{2r-1,2r\}\}.$
    A multigraph $G(\Pi)$ is obtained by identifying the vertices in $\mathcal{G}$ according to the partition $\Pi$. The multigraph $G(\Pi)$ has $r$ edges and $k$ vertices, and each vertex has a degree of at least 2. Also, we focus on the ones that are loop-free, since $A$ is diagonal-free and any loop in $G(\Pi)$ gives a zero in the sum.
  • We further denote $\mathcal{P}_{\geq 2}(2r,k,t)\subset\mathcal{P}_{\geq 2}(2r,k)$ to be the set of partitions of $[2r]$ with exactly $k$ blocks and the corresponding graph $G(\Pi)$ has $t$ connected components. Here we only consider $21\leq t\leq k/$ 2, since we assume $A$ is diagonal-free and each connected component contains at least two vertices.
Using the moment bounds
$\mathbb{E}|x_{i}|^{l}=p_{i}\quad\text{for }l\geq 2,$
and
$\mathbb{E}|X_{i}|^{l}=\mathbb{E}|x_{i}|^{l}\mathbb{E}|\xi_{i}|^{l}\leq p_{i}(K{l})^{l}\leq p_{i}(eK)^{l}l!,$
we can bound (4.5) by
$(eK)^{2r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\left(\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!\sum_{s\in[n]^{\Pi}}\left(\prod_{\{\pi_{1},\pi_{2}\}\in E(G(\Pi))}a_{s_{\pi_{1}}s_{\pi_{2}}}\prod_{\pi\in\Pi}p_{s_{\pi}}\right)\right).$ (4.6)
For simplicity, for a fixed multigraph $G$, we define
$\mathcal{S}_{G}:=\sum_{s\in[n]^{V(G)}}\left(\prod_{\{\pi_{1},\pi_{2}\}\in E(G)}a_{s_{\pi_{1}}s_{\pi_{2}}}\prod_{\pi\in V(G)}p_{s_{\pi}}\right).$ (4.7)
Hence, as $\Pi=V(G(\Pi))$, (4.6) becomes
$\mathbb{E}\left|\sum_{i\neq j}a_{ij}X_{i}X_{j}\right|^{r}$ $\leq(eK)^{2r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\Big{(}\mathcal{S}_{G(\Pi)}\cdot\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!\Big{)}.$ (4.8)
Step 2. (Fix $\Pi$ and bound $\mathcal{S}_{G(\Pi)}$ uniformly). The next lemma provides estimates for $\mathcal{S}_{G(\Pi)}$. Recall that here $\gamma_{1,\infty}:=\max_{i}\left\{\sum_{j\neq i}|a_{ij}|p_{j},\sum_{j\neq i}|a_{ji}|p_{j}\right\}$ for a diagonal-free matrix $A$.
LEMMA 4.2.
For each fixed $\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)$, with $\sigma^{2}\coloneqq\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}$, there is
$\mathcal{S}_{G(\Pi)}$ $\leq\|A\|_{\max}^{r-k}\cdot\gamma_{1,\infty}^{k-2t}\cdot\sigma^{2t}.$ (4.9)
The proof of Lemma 4.2 is deferred to the end of this section. It is worth noting that the bound on $\mathcal{S}_{G(\Pi)}$ in Lemma 4.2 is uniform and independent of $G(\Pi)$. To proceed from (4.8), we now need to estimate
$\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!.$
Step 3. (Consider all possible $\Pi$). For each $G(\Pi)$ with $t$ connected components, we can partition their $r$ edges and $k$ vertices into $t$ components and denote $k_{i},r_{i}$ to be the number of vertices and edges in the $i$-th component correspondingly. More precisely, such a partition $\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)$ can be attained in the following steps:
  1. 1.
    First we consider all the partitions that partition $[2r]$ into $k$ labeled subsets, i.e. each partition is indeed a surjective map $\phi:[2r]\to[k]$. Let $\Phi(\Pi)$ denote the set of such surjection $\phi$ that corresponds to partition $\Pi$. Therefore, as the parts in partition $\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)$ is not ordered, $|\Phi(\Pi)|=k!$. Hence,
    $\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!$ $\leq\frac{1}{k!}\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\sum_{\phi\in\Phi(\Pi)}\prod_{\tau=1}^{k}\deg_{G(\phi)}(\tau)!.$ (4.10)
  2. 2.
    To compute the sum over $\phi$, first consider all possible positive integer series of $(r_{1},\dots,r_{t})$ and $(k_{1},\dots,k_{t})$, such that $\sum_{i=1}^{t}r_{i}=r$ and $\sum_{i=1}^{t}k_{i}=k$. There are $\binom{r-1}{t-1}\binom{k-1}{t-1}\leq 2^{r+k}$ such pairs of series in total.
  3. 3.
    For fixed series $(r_{1},\dots,r_{t})$ and $(k_{1},\dots,k_{t})$, consider the connected components in $G$ are ordered in lexicographic order on the edge partition and the $i$-th component has $r_{i}$ edges and $k_{i}$ vertices. As the order is determined by the edge partition, the edge with the smallest index in the $i$-th component is determined after fixing the edges in the first $i-1$ components. Therefore, there are at most
    $\binom{r-1}{r_{1}-1}\binom{r-r_{1}-1}{r_{2}-1}\cdots\binom{r-r_{1}-\dots-r_{t-1}-1}{r_{t}-1}\leq\ \prod_{i=1}^{t}\binom{r}{r_{i}-1}$
    many ways to partition $r$ edges $\{\{1,2\},\dots,\{2r-1,2r\}\}$ into $t$ connected components such that the $i$-th component has $r_{i}$ edges. Similarly, as we have already ordered the connected components, there are
    $\binom{k}{k_{1}}\binom{k-k_{1}}{k_{2}}\cdots\binom{k-k_{1}-\dots-k_{t-1}}{k_{t}}\leq\ \prod_{i=1}^{t}\binom{k}{k_{i}}$
    many ways to partition the vertices. Therefore,
    $\frac{1}{k!}\sum_{\phi}\prod_{\tau=1}^{k}\deg_{G(\phi)}(\tau)!$ (4.11)
    $\leq\frac{2^{r+k}}{k!}\max_{\text{component partition}}\left(\prod_{i=1}^{t}\binom{r}{r_{i}-1}\binom{k}{k_{i}}\right)\sum_{\phi:\text{consistent with }r_{i},k_{i}}\prod_{\tau=1}^{k}\deg_{G(\phi)}(\tau)!$ (4.12)
    $\leq\frac{2^{r+k}}{k!}\max_{\text{component partition}}\left(\prod_{i=1}^{t}\binom{r}{r_{i}-1}\binom{k}{k_{i}}\right)\left(\prod_{i=1}^{t}\sum_{\phi_{i}:[2r_{i}]\to[k_{i}]}\prod_{\tau=1}^{k_{i}}\deg_{G(\phi_{i})}(\tau)!\right).$ (4.13)
  4. 4.
    Finally, inside $i$-th connected component, there are $r_{i}$ edges and $k_{i}$ labeled vertices $\{j_{1},\dots,j_{k_{i}}\}$. Let us further fix the degree $d_{1},\dots,d_{k_{i}}$ of each vertex and then sum over all possible degree sequences. As the vertices are labeled and the degree of each vertex is fixed, there are at most $\binom{2r_{i}}{d_{1},\dots,d_{k_{i}}}$ many different partition $\phi_{i}$ of $[2r]$. Therefore, with $d_{1}+\dots+d_{k_{i}}=2r_{i}$, we have
    $\sum_{\phi_{i}:[2r_{i}]\to[k_{i}]}\prod_{\tau=1}^{k}\deg_{G(\phi_{i})}(\tau)!$ $\leq\sum_{d_{1},\dots,d_{k_{i}}}\binom{2r_{i}}{d_{1}+\dots+d_{k_{i}}=2r_{i}}\cdot\prod_{\ell=1}^{k_{i}}d_{\ell}!$ (4.14)
    $\leq\binom{2r_{i}-1}{k_{i}-1}(2r_{i})!$ (4.15)
    $\leq 2^{4r_{i}}r_{i}^{2r_{i}}.$ (4.16)
Therefore, combining the steps above, with $r_{i}\geq 2$ we have
$\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!$ $\leq\frac{2^{r+k}}{k!}\max_{\text{component partition}}\left(\prod_{i=1}^{t}\binom{r}{r_{i}-1}\binom{k}{k_{i}}\right)\left(\prod_{i=1}^{t}2^{4r_{i}}r_{i}^{2r_{i}}\right)$
$\leq\frac{2^{r+k}}{k!}\max_{\sum r_{i}=r,\sum k_{i}=k}\prod_{i=1}^{t}\left(\frac{er}{r_{i}-1}\right)^{r_{i}-1}\left(\frac{ek}{k_{i}}\right)^{k_{i}}2^{4r_{i}}r_{i}^{2r_{i}}$
$\leq\frac{C^{r}}{k!}\max_{\sum r_{i}=r,\sum k_{i}=k}\prod_{i=1}^{t}\left(\frac{r}{r_{i}}\right)^{r_{i}-1}\left(\frac{k}{k_{i}}\right)^{k_{i}}r_{i}^{2r_{i}}$
$\leq\frac{C^{r}r^{2r}}{k!}\max_{\sum r_{i}=r,\sum k_{i}=k}\prod_{i=1}^{t}\left(\frac{r}{r_{i}}\right)^{r_{i}-1}\left(\frac{k}{k_{i}}\right)^{k_{i}}\left(\frac{r_{i}}{r}\right)^{2r_{i}}$
$\underset{(a)}{\leq}\frac{C^{r}r^{2r}}{k!}\max_{\sum r_{i}=r,\sum k_{i}=k}\prod_{i=1}^{t}\left(\frac{r_{i}}{r}\right)^{r_{i}-k_{i}+1}.$ (4.17)
In the inequality (a), we applied Gibbs’ Inequality to the $(k/k_{i})^{k_{i}}$ terms, which is introduced in Lemma A.1.
To proceed from (4.1), we again apply Gibbs’ inequality. Denote $z_{i}\coloneqq r_{i}-k_{i}+1\geq 1$ and $Z=\sum_{i=1}^{t}z_{i}=r-k+t$. We then have
$\prod_{i=1}^{t}\left(\frac{r_{i}}{r}\right)^{z_{i}}\leq\prod_{i=1}^{t}\left(\frac{z_{i}}{Z}\right)^{z_{i}}=Z^{-Z}\prod_{i=1}^{t}z_{i}^{z_{i}}\leq Z^{-Z}(Z-t+1)^{Z-t+1}.$ (4.18)
Here, the last inequality follows from Lemma A.2 and that the maximum of $\prod_{i}z_{i}^{z_{i}}$ is attained when $t-1$ of the $z_{i}$’s are equal to 1, and the remaining $z_{i}$ equals $Z-(t-1)$. Therefore, we have
$\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!$ $\leq\frac{C^{r}r^{2r}}{k!}\cdot\frac{(r-k+1)^{r-k+1}}{(r-k+t)^{r-k+t}}$ (4.19)
$\leq C^{r}r^{2r}k^{-k}{t^{-(t-1)}}$ (4.20)
$\leq C^{r}r^{2r-k-t+1}\left(\frac{r}{k}\right)^{k}\left(\frac{r}{t}\right)^{t}$ (4.21)
$\leq C^{r}r^{2r-k-t}.$ (4.22)
Here in the last inequality, we use the fact that $f(x)=(r/x)^{x}$ attains its maximum at $x=r/e$.
Step 4. (Conclusion). For the last part of the proof, combining the results from Lemma 4.2, rearranging terms, and applying Young’s inequality, we obtain the following moment bound. Here $\sigma^{2}\coloneqq\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}$. Continuing from (4.8), we have
$\mathbb{E}\left|\sum_{i\neq j}a_{ij}X_{i}X_{j}\right|^{r}$ $\leq(eK)^{2r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\Big{(}\mathcal{S}_{G(\Pi)}\cdot\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!\Big{)}$ (4.23)
$\leq(eK)^{2r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\|A\|_{\max}^{r-k}\gamma_{1,\infty}^{k-2t}\sigma^{2t}\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!$ (4.24)
$\leq(CK)^{2r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\|A\|_{\max}^{r-k}\gamma_{1,\infty}^{k-2t}\sigma^{2t}\cdot r^{2r-k-t}$ (4.25)
$\leq C^{2r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}(r^{2}K^{2}\|A\|_{\max})^{r-k}(rK^{2}\gamma_{1,\infty})^{k-2t}(\sqrt{r}K^{2}\sigma)^{2t}$ (4.26)
$\leq C^{2r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\max\left\{\frac{(r^{2}K^{2}\|A\|_{\max})^{r}}{r/(r-k)},\frac{(rK^{2}\gamma_{1,\infty})^{r}}{r/(k-2t)},\frac{(\sqrt{r}K^{2}\sigma)^{r}}{r/2t}\right\}$ (4.27)
$\leq C^{2r}r^{2}\max\left\{{(r^{2}K^{2}\|A\|_{\max})^{r}},{(rK^{2}\gamma_{1,\infty})^{r}},{(\sqrt{r}K^{2}\sigma)^{r}}\right\}$ (4.28)
$\leq C_{1}^{r}\max\left\{{(r^{2}K^{2}\|A\|_{\max})^{r}},{(rK^{2}\gamma_{1,\infty})^{r}},{(\sqrt{r}K^{2}\sigma)^{r}}\right\},$ (4.29)
where $C,C_{1}$ are absolute constants.
4.2 Proof of Lemma 4.2
For each fixed $\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)$ and its corresponding graph $G:=G(\Pi)$, we aim to bound
$\mathcal{S}_{G}:=\sum_{s\in[n]^{V(G)}}\left(\prod_{\{\pi_{1},\pi_{2}\}\in E(G)}a_{s_{\pi_{1}}s_{\pi_{2}}}\cdot\prod_{\pi\in V(G)}p_{s_{\pi}}\right).$ (4.30)
using the matrix norms of $A$. Recall that the graph $G$ contains $r$ edges, and each vertex has a degree of at least 2. For diagonal-free $A$, we recall that
$\gamma_{1,\infty}:=\max_{i}\left\{\sum_{j\neq i}|a_{ij}|p_{j},\sum_{j\neq i}|a_{ji}|p_{j}\right\},\quad\sigma^{2}\coloneqq\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}.$
The main idea for the estimates is to delete the edges in $G=G(\Pi)$ as the following steps:
  1. 1.
    For each connected component $\mathcal{C}$ of $G$, we identify a spanning tree $T$ and add one arbitrarily chosen edge from $E(G)\setminus E(T)$, creating a cycle. We then delete all other edges in $\mathcal{C}$. Let $G^{\prime}$ be the remaining subgraph after deletion to all connected components, and hence $|E(G^{\prime})|=k$.
    For each deleted edge, say $\{u,v\}$, we bound the corresponding $a_{uv}\leq\|A\|_{\max}$, and we will delete $r-k$ many edges this way in total, which implies that we delete a total of $r-k$ edges from $G$. These deleted edges collectively contribute a factor of $\|A\|_{\max}^{r-k}$ to bound $\mathcal{S}_{G}$, i.e.
    $\mathcal{S}_{G}\leq\|A\|_{\max}^{r-k}\cdot\mathcal{S}_{G^{\prime}}$
  2. 2.
    Now $G^{\prime}$ is a graph consisting of $t$ connected components, each containing exactly one cycle and possibly some disjoint branches (or sub-trees) connected to the cycles. Let $G^{\prime\prime}\subseteq G^{\prime}$ denote the smaller subgraph containing all $t$ cycles, denoted by $C_{j}$ for $j=1,\ldots,t$. From $G^{\prime}$ to $G^{\prime\prime}$, we can repeatedly remove vertex $v$ and its edge $\{u,v\}$, where the degree of $v$ is 1. We fix $u$ first and bound $\sum_{v=1}^{n}a_{uv}p_{v}\leq\gamma_{1,\infty}$ or $\sum_{v=1}^{n}a_{vu}p_{v}\leq\gamma_{1,\infty}.$ Suppose $\ell$ edges are removed in this step, we have $|E(G^{\prime\prime})|=k-\ell$, and more importantly,
    $\mathcal{S}_{G^{\prime}}\leq\gamma_{1,\infty}^{\ell}\cdot\mathcal{S}_{G^{\prime\prime}}.$
    Moreover, $\mathcal{S}_{G^{\prime\prime}}=\prod_{j=1}^{t}\mathcal{S}_{C_{j}}.$
  3. 3.
    Finally, for each remaining cycle $C_{j}$ denoted by $(v_{1},v_{2},\cdots,v_{m})$ with $m$ edges, we have
    $\sum_{\begin{subarray}{c}s\in[n]^{V(C_{j})}\end{subarray}}a_{s(v_{1})s(v_{2})}a_{s(v_{2})s(v_{3})}\cdots a_{s(v_{m})s(v_{1})}\cdot p_{s(v_{1})}p_{s(v_{2})}\cdots p_{s(v_{m})}$ (4.31)
    $\quad=\sum_{\begin{subarray}{c}i_{1},\dots,i_{m}=1\end{subarray}}^{n}a_{i_{1}i_{2}}a_{i_{2}i_{3}}\cdots a_{i_{m}i_{1}}\cdot p_{i_{1}}p_{i_{2}}\cdots p_{i_{m}}$ (4.32)
    $\quad=\operatorname{tr}(B^{m})\leq\|B\|_{F}^{2}\|B\|^{m-2},$ (4.33)
    where $B:=\sqrt{P}A\sqrt{P}$ and $P:=\mathrm{diag}(p_{1},\dots,p_{n})$. Next, we show that $\|B\|\leq\gamma_{1,\infty}$. To see this, assume that $\|B\|$ is attained at some vector $x=(x_{1},\dots,x_{n})^{T}\in\mathbb{S}^{n-1}$, i.e., $\|B\|=\|Bx\|$. By Cauchy-Schwartz inequality, we have
    $\|B\|^{2}$ $=\|Bx\|^{2}=\sum_{i}\left(\sum_{j}a_{ij}\sqrt{p_{i}p_{j}}x_{j}\right)^{2}\leq\sum_{i}\left(\sum_{j}a_{ij}{p_{j}}\right)\left(\sum_{j}a_{ij}{p_{i}}x_{j}^{2}\right)$ (4.34)
    $\leq\gamma_{1,\infty}\sum_{i}\left(\sum_{j}a_{ij}{p_{i}}x_{j}^{2}\right)=\gamma_{1,\infty}\sum_{j}\left(\sum_{i}a_{ij}{p_{i}}x_{j}^{2}\right)$ (4.35)
    $=\gamma_{1,\infty}^{2}\sum_{j}x_{j}^{2}=\gamma_{1,\infty}^{2}.$ (4.36)
    It follows that $\operatorname{tr}(B^{m})\leq\|B\|_{F}^{2}\|B\|^{m-2}\leq\sigma^{2}\gamma_{1,\infty}^{m-2}$, where $\|B\|_{F}^{2}=\sum_{i,j}a_{ij}^{2}p_{i}p_{j}=\sigma^{2}$ is the variance of the quadratic form. Hence,
    $\mathcal{S}_{G^{\prime\prime}}=\prod_{j=1}^{t}\mathcal{S}_{C_{j}}\leq\sigma^{2t}\gamma_{1,\infty}^{k-\ell-2t}.$
In the process above, we have $r-k$ edges deleted in the first step and $\ell$ edges deleted in the second step. Then
$\mathcal{S}_{G}$ $\leq\|A\|_{\max}^{r-k}\cdot\mathcal{S}_{G^{\prime}}\leq\|A\|_{\max}^{r-k}\gamma_{1,\infty}^{\ell}\cdot\mathcal{S}_{G^{\prime\prime}}\leq\|A\|_{\max}^{r-k}\gamma_{1,\infty}^{k-2t}\sigma^{2t}.$ (4.37)
4.3 Proof of Theorem 2
The proof of Theorem 2 follows a similar moment estimation scheme to that of Theorem 1. We provide a brief proof here, focusing on the differences. First of all, the moments of $\alpha$-sub-exponential random variables satisfy
$\mathbb{E}|\xi_{i}|^{l}\leq K^{l}l^{\frac{l}{\alpha}}.$
For $\alpha\geq 1$, we simply bound $\mathbb{E}|\xi_{i}|^{l}\leq K^{l}l^{l}$, which leads to results identical to those of Theorem 1. We will focus on the case where $\alpha\in(0,1)$. Since
$\mathbb{E}|X_{i}|^{l}=\mathbb{E}|x_{i}|^{l}\mathbb{E}|\xi_{i}|^{l}\leq p_{i}(eK)^{l}l!\cdot l^{(\frac{1}{\alpha}-1)l},$
analogous to the derivation of (4.6), we now have
$\mathbb{E}|X^{T}AX|^{r}\leq(eK)^{2r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\sum_{\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)}\left(\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)!\deg_{G(\Pi)}(\tau)^{(\frac{1}{\alpha}-1)\deg_{G(\Pi)}(\tau)}\mathcal{S}_{G(\Pi)}\right),$ (4.38)
where $\mathcal{P}_{\geq 2}(2r,k,t)$ and the definitions of $G(\Pi),\mathcal{S}_{G(\Pi)}$ follow those in (4.6).
Note that compared to (4.6), we only have an extra multiplicative factor
$\left(\prod_{\tau\in\Pi}\deg_{G(\Pi)}(\tau)^{\deg_{G(\Pi)}(\tau)}\right)^{\frac{1}{\alpha}-1}:=\left(\mathcal{D}(\Pi)\right)^{\frac{1}{\alpha}-1}.$
We claim that $\mathcal{D}(\Pi)\leq 2^{2r}r^{2(r-k+1)}$ uniformly for all $\Pi\in\mathcal{P}_{\geq 2}(2r,k,t)$. Consider any fixed graph $G(\Pi)$ with $r$ edges and $k$ vertices. Let its degree sequence be denoted by $d_{1},\dots,d_{k}$. We have $d_{i}\geq 2$ and $d_{1}+\dots+d_{k}=2r$. Therefore, by Lemma A.2
$\prod_{i=1}^{k}d_{i}^{d_{i}}\leq 2^{2}\times\dots\times 2^{2}\times(2r-2(k-1))^{2r-2(k-1)}=2^{2r}r^{2(r-k+1)}.$ (4.39)
Using the estimates on this additional factor $\left(\mathcal{D}(\Pi)\right)^{\frac{1}{\alpha}-1}$ and then following a similar estimation procedure as in the proof of Lemma 4.1, we obtain
$\mathbb{E}|X^{T}AX|^{r}$ $\leq(CK^{2})^{r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\|A\|_{\max}^{r-k}\gamma_{1,\infty}^{k-2t}\sigma^{2t}r^{2r-k-t}\cdot\big{(}2^{2r}r^{2(r-k+1)}\big{)}^{\frac{1}{\alpha}-1}$ (4.40)
$\leq(C_{\alpha}K^{2})^{r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\|A\|_{\max}^{r-k}\gamma_{1,\infty}^{k-2t}\sigma^{2t}r^{2r-k-t}\cdot r^{2(\frac{1}{\alpha}-1)(r-k+1)}$ (4.41)
$\leq(C_{\alpha}K^{2})^{r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}\|A\|_{\max}^{r-k}\gamma_{1,\infty}^{k-2t}\sigma^{2t}\cdot r^{\frac{2}{\alpha}r-(\frac{2}{\alpha}-1)k-t}$ (4.42)
$=(C_{\alpha}K^{2})^{r}\sum_{k=2}^{r}\sum_{t=1}^{k/2}(r^{\frac{2}{\alpha}}\|A\|_{\max})^{r-k}(r\gamma_{1,\infty})^{k-2t}(\sqrt{r}\sigma)^{2t}$ (4.43)
$\leq(C_{\alpha}K^{2})^{r}\max\left\{r^{\frac{2}{\alpha}}\|A\|_{\max},r\gamma_{1,\infty},\sqrt{r}\sigma\right\}^{r}.$ (4.44)
Here in the last inequality, we use Young’s inequality for the products of three terms.
Hence,
$\|X^{T}AX\|_{L_{r}}\leq C_{\alpha}K^{2}\max\left\{r^{\frac{2}{\alpha}}\|A\|_{\max},r\gamma_{1,\infty},\sqrt{r}\sigma\right\}$
for any even integer $r$.
From the above norm bound, by similar discussion in the proof to Theorem 1, we derive the tail inequality. By Markov’s inequality,
$\mathbb{P}\left(\left|\sum_{i\not=j}a_{ij}X_{i}X_{j}\right|\geq t\right)\leq\left(\frac{1}{t}C_{\alpha}K^{2}\max\left\{r^{\frac{2}{\alpha}}\|A\|_{\max},r\gamma_{1,\infty},\sqrt{r}\sigma\right\}\right)^{r}.$ (4.45)
We choose $r$ to be the largest even integer such that
$2\leq r\leq\min\left\{\frac{t^{2}}{C_{\alpha}e^{2}K^{4}\sigma^{2}},\frac{t}{C_{\alpha}eK^{2}\gamma_{1,\infty}},\left(\frac{t}{C_{\alpha}eK^{2}\|A\|_{\max}}\right)^{\frac{\alpha}{2}}\right\}$ (4.46)
and this implies
$\frac{1}{t}C_{\alpha}K^{2}\max\left\{r^{\frac{2}{\alpha}}\|A\|_{\max},r\gamma_{1,\infty},\sqrt{r}\sigma\right\}\leq\frac{1}{e}.$ (4.47)
If the right-hand side of (4.46) is less than 2, then we choose $C^{\prime}=e^{2}$ such that the conclusion of Theorem 2 trivially holds. This proves Theorem 2.
5 Proofs of Theorem 3 and Corollary 1
Following the proof of the off-diagonal case, we now focus on the concentration of the linear form. Note that given Theorem 3, Corollary 1 immediately holds as
$\mathbb{E}|X_{i}^{2}-\mathbb{E}X_{i}^{2}|^{l}\leq 2^{l}\mathbb{E}|X_{i}|^{2l}\leq 2^{l}p_{i}K^{2l}(2l)^{\frac{2l}{\alpha}}.$ (5.1)
Here in the first inequality, we use the fact that for a non-negative random variable $Z$, one has
$\mathbb{E}(Z-\mathbb{E}Z)^{l}\leq\mathbb{E}[2^{l-1}(Z^{l}+(\mathbb{E}Z)^{l})]\leq 2^{l}\mathbb{E}Z^{l}.$
In the rest of this section, we will prove Theorem 3 by moment method. The following lemma provide a moment bound for the linear sum.
LEMMA 5.1.
With random vector $X$ defined in Theorem 3 and any positive even number $r$, we have the moment bound
$\mathbb{E}\left|\sum_{i=1}^{n}a_{i}X_{i}\right|^{r}$ $\leq(C_{\alpha}K^{2})^{r}\max\left\{r\|a\|_{\infty},\sqrt{r\sum_{i=1}^{n}a_{i}^{2}p_{i}}\right\}^{r}$
Proof.
Note that for a single variable we have $\mathbb{E}|X_{i}|^{l}\leq p_{l}K^{l}l^{\frac{l}{\alpha}}$. Let $\mathcal{P}(r,k)$ denote the set of all the partitions of $[r]$ into $k$ subsets and $\mathcal{P}_{\geq 2}(r,k)$ denote the subset of those partitions where all blocks have size at least 2. Furthermore, let $\lambda\coloneqq\sqrt{\sum_{i=1}^{n}a_{i}^{2}p_{i}}$. Since the $Y_{i}$’s are centered and satisfy the moment bounds in (5.1), we then obtain the following estimates:
$\mathbb{E}\left|\sum_{i=1}^{n}a_{i}X_{i}\right|^{r}$ $\leq\sum_{i_{1},\dots,i_{r}}a_{i_{1}}\dots a_{i_{r}}\mathbb{E}[X_{i_{1}}\dots X_{i_{r}}]$ (5.2)
$\leq\sum_{k=1}^{r/2}\sum_{\Pi\in\mathcal{P}_{\geq 2}(r,k)}\sum_{s\in[n]^{\Pi}}\prod_{\tau\in\Pi}a_{s(\tau)}^{|\tau|}\mathbb{E}|X_{s(\tau)}|^{|\tau|}$ (5.3)
$\leq\sum_{k=1}^{r/2}\sum_{\Pi\in\mathcal{P}_{\geq 2}(r,k)}\prod_{\tau\in\Pi}\left(\sum_{s=1}^{n}a_{s}^{|\tau|}\mathbb{E}|X_{s}|^{|\tau|}\right)$ (5.4)
$\underset{(a)}{\leq}\sum_{k=1}^{r/2}\sum_{\Pi\in\mathcal{P}_{\geq 2}(r,k)}\prod_{\tau\in\Pi}\left(K^{|\tau|}|\tau|^{\frac{|\tau|}{\alpha}}\sum_{s=1}^{n}\|a\|_{\infty}^{|\tau|-2}a_{s}^{2}p_{s}\right)$ (5.5)
$=K^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\sum_{\Pi\in\mathcal{P}_{\geq 2}(r,k)}\left(\prod_{\tau\in\Pi}|\tau|^{|\tau|}\right)^{\frac{1}{\alpha}}$ (5.6)
$\underset{(b)}{\leq}K^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\frac{1}{k!}\sum_{\Pi\in\mathcal{P}_{\geq 2}(r,k)}\sum_{\phi\in\Phi(\Pi)}\left(\prod_{i=1}^{k}|\phi^{-1}(i)|^{|\phi^{-1}(i)|}\right)^{\frac{1}{\alpha}}$ (5.7)
=$\underset{(c)}{=}K^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\frac{1}{k!}\sum_{\begin{subarray}{c}\phi:[r]\to[k],\\ |\phi^{-1}(i)|\geq 2\end{subarray}}\left(\prod_{i=1}^{k}|\phi^{-1}(i)|^{|\phi^{-1}(i)|}\right)^{\frac{1}{\alpha}}$ (5.8)
$=K^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\frac{1}{k!}\sum_{r_{1}+\dots+r_{k}=r;r_{i}\geq 2}\binom{r}{r_{1},\dots,r_{k}}\left(\prod_{i=1}^{k}r_{i}^{r_{i}}\right)^{\frac{1}{\alpha}}$ (5.9)
$\leq(C_{\alpha}K)^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\frac{r^{r}}{k^{k}}\sum_{r_{1}+\dots+r_{k}=r;r_{i}\geq 2}\left(\prod_{i=1}^{k}r_{i}^{r_{i}}\right)^{\frac{1}{\alpha}-1}.$ (5.10)
In the inequality (a) above, we use the the fact that $|\tau|\geq 2$ for $\tau\in\Pi$ so $a_{s}^{|\tau|-2}\leq\|a\|_{\infty}^{|\tau|-2}$. In the inequality (b), similar to the proof of Lemma 4.2, we use $\Phi(\Pi)$ to denote the set of surjections $\phi:[r]\to[k]$ corresponding to the partition $\Pi$, and further in (c) we rewrite the sum over $\Pi$ to surjections $\phi$.
Depending on the value of $\alpha$, we have the following two cases:
  • When $\frac{1}{\alpha}\geq 1$, applying Lemma A.2, for any fixed series $r_{1},\dots,r_{k}$, we have
    $\prod_{i=1}^{k}r_{i}^{r_{i}}\leq 2^{2(k-1)}(r-2(k-1))^{r-2(k-1)}\leq 4^{k}r^{r-2k+2}.$
    Therefore, we further obtain
    $\mathbb{E}\left|\sum_{i=1}^{n}a_{i}X_{i}\right|^{r}$ $\leq(C_{\alpha}K)^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot\frac{r^{r}}{k^{k}}\sum_{r_{1}+\dots+r_{k}=r;r_{i}\geq 2}\left(\prod_{i=1}^{k}r_{i}^{r_{i}}\right)^{\frac{1}{\alpha}-1}$ (5.11)
    $\leq(C_{\alpha}K)^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot\frac{r^{r}}{k^{k}}\binom{r-1}{k-1}\left(4^{k}r^{r-2k+2}\right)^{\frac{1}{\alpha}-1}$ (5.12)
    $\leq(C_{\alpha}K^{2})^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot r^{\frac{1}{\alpha}r-(\frac{2}{\alpha}+1)k}$ (5.13)
    $\leq(C_{\alpha}K)^{r}\sum_{k=1}^{r/2}(r^{\frac{1}{\alpha}}\|a\|_{\infty})^{r-2k}(\sqrt{r}\lambda)^{2k}$ (5.14)
    $\underset{(d)}{\leq}(C_{\alpha}K)^{r}\max_{1\leq k\leq r/2}\max\left\{\frac{(r^{\frac{1}{\alpha}}\|a\|_{\infty})^{r}}{r/(r-2k)},\frac{(\sqrt{r}\lambda)^{r}}{r/(2k)}\right\}$ (5.15)
    $\leq(C_{\alpha}K)^{r}\max\left\{r^{\frac{1}{\alpha}}\|a\|_{\infty},\sqrt{r}\lambda\right\}^{r},$ (5.16)
    where the inequality (d) is due to Young’s inequality.
  • When $\frac{1}{\alpha}<1$, simply bounding the term of $(\prod_{i=1}^{k}r_{i}^{r_{i}})^{\frac{1}{\alpha}-1}$ by 1 and using the Young’s inequality, we have
    $\mathbb{E}\left|\sum_{i=1}^{n}a_{i}X_{i}\right|^{r}$ $\leq(C_{\alpha}K^{2})^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot\frac{r^{r}}{k^{k}}\sum_{r_{1}+\dots+r_{k}=r;r_{i}\geq 2}\left(\prod_{i=1}^{k}r_{i}^{r_{i}}\right)^{\frac{1}{\alpha}-1}$
    $\leq(C_{\alpha}K^{2})^{r}\sum_{k=1}^{r/2}\|a\|_{\infty}^{r-2k}\lambda^{2k}\cdot{r^{r-k}}\binom{r-1}{k-1}$
    $\leq(C_{\alpha}K^{2})^{r}\max\left\{r\|a\|_{\infty},\sqrt{r}\lambda\right\}^{r}$ (5.17)
Finally, Theorem 3 can be proved similarly to the last part of the proof of Theorem 1. With $C^{\prime}=e^{2}$, either the right-hand side of the inequality is at least one, or we can set $r$ to be the largest positive even integer satisfying
$2\leq r\leq\min\left\{\left(\frac{t}{C_{\alpha}eK^{2}\|a\|_{\infty}}\right)^{\min\{\alpha,1\}},\frac{t^{2}}{C_{\alpha}^{2}e^{2}K^{4}\lambda^{2}}\right\},$
completing the proof of Theorem 3.
6 Proofs of applications
6.1 Proof of Theorem 8
In this section, we present the proof of Theorem 8. While our proof strategy closely follows the framework developed in [24] (and references therein), we focus on highlighting the key differences.
The new technical ingredients in our proof are the large deviation inequalities for sparse $\alpha$-subexponential random variables. We first note that the large deviation inequalities in Theorem 2 and Corollary 1 can be extended to complex coefficients or complex matrices by treating the real and imaginary parts separately. For the convenience of our subsequent analysis, we summarize the necessary results below.
Let $a=(a_{1},\cdots,a_{n})^{T}\in\mathbb{C}^{n}$ and $A=(a_{ij})_{1\leq i,j\leq n}\in\mathbb{C}^{n\times n}$, where $A$ is diagonal-free and $A=A^{*}$. Consider a random vector $X=(X_{1},\cdots,X_{n})^{T}\in\mathbb{R}^{n}$ with independent centered $\alpha$-subexponential random components. Each $X_{i}=x_{i}\xi_{i}$, where $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p)$ and $\xi_{i}$ has mean 0, variance 1, and satisfies $\|\xi_{i}\|_{L_{r}}\leq Kr^{\frac{1}{\alpha}}$ for all $r\geq 1$. Then by Theorem 2, Theorem 3, and Corollary 1, there is a constant $C_{\alpha}>0$ such that for any $r\geq 2$,
$\left\|\sum_{i=1}^{n}a_{i}X_{i}\right\|_{L_{r}}\leq C_{\alpha}K\max\left\{r^{\max\{\frac{1}{\alpha},1\}}\|a\|_{\infty},\sqrt{r}\sqrt{p}\|a\|_{2}\right\},$ (6.1)
$\left\|\sum_{i=1}^{n}a_{i}(X_{i}^{2}-p)\right\|_{L_{r}}\leq C_{\alpha}K^{2}\max\left\{r^{\max\{\frac{2}{\alpha},1\}}\|a\|_{\infty},\sqrt{r}\sqrt{p}\|a\|_{2}\right\},$ (6.2)
$\|X^{T}AX\|_{L_{r}}\leq C_{\alpha}K^{2}\max\left\{r^{\max\{\frac{2}{\alpha},2\}}\|A\|_{\max},rp\|A\|_{1,\infty},\sqrt{r}p\|A\|_{F}\right\}.$ (6.3)
Without loss of generality, we assume $C_{\alpha}\geq 1$.
In order to prove Theorem 8, it suffices to prove the following result. Define a fundamental error parameter
$\zeta\equiv\zeta(n,r,p,\eta,\alpha):=\sqrt{12C_{\alpha}K^{2}\max\left\{\frac{r^{\max\{\frac{1}{\alpha},1\}}}{\sqrt{np}},\frac{r}{\sqrt{n\eta}}\right\}}.$ (6.4)
THEOREM 12.
Let $z\in\mathbf{S}$, $r\in\mathbb{N}$ and $p\in(0,1)$. Assume $n\geq 10$, $r\geq 2$ and $r^{\max\{\frac{1}{\alpha},1\}}\leq\sqrt{np}$. For any $t\geq 1$ such that $20t\zeta\leq 1$, we have
$\mathbb{P}\left(\max_{i,j}|G_{ij}(z)-m(z)\delta_{ij}|>12t\zeta\right)\leq 3n^{6}t^{-2r}.$
Let us assume Theorem 12 holds and prove Theorem 8. Given $D>0$ and $\delta\in(0,1)$, we take $r=\log n$, $t=\exp(\frac{D+7}{2})$. Let $\mathtt{f}(D,\delta):=\max\{1,\frac{(20\delta)^{4}}{12^{4}}\}.$ Then the conditions of Theorem 12 are satisfied provided that
$p\geq\frac{12^{6}e^{2(D+7)}\mathtt{f}(D,\delta)}{\delta^{4}}\cdot\frac{C_{\alpha}^{2}K^{4}(\log n)^{\max\{\frac{2}{\alpha},2\}}}{n}$
and
$\eta\geq\frac{12^{6}e^{2(D+7)}\mathtt{f}(D,\delta)}{\delta^{4}}\cdot\frac{C_{\alpha}^{2}K^{4}(\log n)^{2}}{n}.$
Under these conditions, we have $20t\zeta\leq 1$, $12t\zeta\leq\delta$ and $3n^{6}t^{-2r}\leq n^{-D}$. Setting
$L(\delta,D):=\frac{12^{6}e^{2(D+7)}\mathtt{f}(D,\delta)}{\delta^{4}},$ (6.5)
we complete the proof of Theorem 8.
The remainder of this section is devoted to the proof of Theorem 12. We begin by introducing necessary notations and preliminary results. Following the notations in [24], for any $k\in[n]$, we denote by
$H^{(k)}:=(H_{ij})_{i,j\in[n]\setminus\{k\}}$
the $(n-1)\times(n-1)$ matrix obtained by deleting the $k$-th row and column from $H$. The Green function of $H^{(k)}$ is defined as
$G^{(k)}(z)=(H^{(k)}-z)^{-1}.$
Moreover, we use the shorthand notation $\sum_{i}^{(k)}:=\sum_{i:i\neq k}$. The Stieltjes transform of the ESD $\mu_{n}$ is given by
$s(z):=\frac{1}{n}\sum_{i=1}^{n}G_{ii}(z).$
Define the random $z$-dependent error parameter
$\Gamma(z):=\max\left\{\max_{i,j}|G_{ij}(z)|,\max_{i,j\neq k}|G_{ij}^{(k)}(z)|\right\}$
and the $z$-dependent indicator function
$\phi(z):=\mathbf{1}(\Gamma(z)\leq 2).$
For brevity, we omit the spectral parameter $z$ when the context is clear. Using Schur’s complement (see Lemma 5.2 from [10]), one has
$\frac{1}{G_{ii}}=-z-s+Y_{i},$ (6.6)
where
$Y_{i}:=H_{ii}+\frac{1}{n}\sum_{k}\frac{G_{ki}G_{ik}}{G_{ii}}-\sum_{k\neq l}^{(i)}H_{ik}G_{kl}^{(i)}H_{li}-\sum_{k}^{(i)}(H_{ik}^{2}-n^{-1})G_{kk}^{(i)}.$ (6.7)
An important identity, known as the Ward identity, is frequently used in our estimates (see (4.4) from [24]):
$\sum_{j}|G_{ij}|^{2}=\frac{\operatorname{Im}G_{ii}}{\eta}.$ (6.8)
The proof of Theorem 12 follows from a standard stochastic continuity argument (see [24, 10]) combined with a bootstrapping technique, proceeding from large scales ($\eta=1$) down to small scales ($\eta\approx\log n/n$). The new technical inputs are the following error estimates under the event $\phi=1$, which are proved using (6.1), (6.2) and (6.3).
LEMMA 6.1 (Main estimates).
Assume $r\geq 2$ and $r^{\max\{\frac{1}{\alpha},1\}}\leq\sqrt{np}$. Then we have
$\max_{i\neq j}\|\phi G_{ij}\|_{L_{r}}\leq 4C_{\alpha}K\max\left\{\frac{r^{\max\{\frac{1}{\alpha},1\}}}{\sqrt{np}},\frac{\sqrt{r}}{\sqrt{n\eta}}\right\},$ (6.9)
$\max_{i,j\neq k}\|\phi(G_{ij}-G_{ij}^{(k)})\|_{L_{r}}\leq 4C_{\alpha}K\max\left\{\frac{r^{\max\{\frac{1}{\alpha},1\}}}{\sqrt{np}},\frac{\sqrt{r}}{\sqrt{n\eta}}\right\},$ (6.10)
$\max_{i}\|\phi Y_{i}G_{ii}\|_{L_{r}}\leq 12C_{\alpha}K^{2}\max\left\{\frac{r^{\max\{\frac{1}{\alpha},1\}}}{\sqrt{np}},\frac{r}{\sqrt{n\eta}}\right\}.$ (6.11)
Proof of Lemma 6.1.
We define
$\phi^{(i)}:=\mathbf{1}(\max_{k,l\neq i}|G_{kl}^{(i)}|\leq 2)$
and observe that $\phi\leq\phi^{(i)}$. Due to the independence of $H^{(i)}$ and the $i$-th row and column of $H$, let $\|\cdot\|_{L_{r}|H^{(i)}}$ denote the conditional $L_{r}$ norm with respect to the conditional expectation $\mathbb{E}(\cdot|H^{(i)})$.
We begin with the proof of (6.9). By Schur’s complement formula (see also Lemma 4.2 from [24]), for $i\neq j$, we have
$G_{ij}=-G_{ii}\sum_{k}^{(i)}H_{ik}G_{kj}^{(i)}.$
Using the nesting property of conditional $L_{r}$ norms, we obtain
$\|\phi G_{ij}\|_{L_{r}}\leq 2\left\|\phi^{(i)}\Big{\|}\sum_{k}^{(i)}H_{ik}G_{kj}^{(i)}\Big{\|}_{L_{r}|H^{(i)}}\right\|_{L_{r}}.$
By (6.1), we derive
$\phi^{(i)}\Big{\|}\sum_{k}^{(i)}H_{ik}G_{kj}^{(i)}\Big{\|}_{L_{r}|H^{(i)}}$ $=\frac{1}{\sqrt{np}}\phi^{(i)}\Big{\|}\sum_{k}^{(i)}\widetilde{H}_{ik}G_{kj}^{(i)}\Big{\|}_{L_{r}|H^{(i)}}$ (6.12)
$\leq\frac{C_{\alpha}K}{\sqrt{np}}\phi^{(i)}\max\left\{r^{\max\{\frac{1}{\alpha},1\}}\max_{k\neq i}|G_{kj}^{(i)}|,\sqrt{rp}\sqrt{\sum_{k}^{(i)}|G_{kj}^{(i)}|^{2}}\right\}$ (6.13)
$\leq 2\frac{C_{\alpha}K}{\sqrt{np}}\max\left\{r^{\max\{\frac{1}{\alpha},1\}},\frac{\sqrt{rp}}{\sqrt{\eta}}\right\},$ (6.14)
where in the last inequality, we used (6.8) to get $\sum_{k}^{(i)}|G_{kj}^{(i)}|^{2}={\operatorname{Im}G_{jj}^{(i)}}/{\eta}.$ It follows that
$\|\phi G_{ij}\|_{L_{r}}\leq 4C_{\alpha}K\max\left\{\frac{r^{\max\{\frac{1}{\alpha},1\}}}{\sqrt{np}},\frac{\sqrt{r}}{\sqrt{n\eta}}\right\}$
and hence (6.9) is proved.
The proof of (6.10) follows analogously by using the identity (see Lemma 4.2 in [24])
$\|\phi(G_{ij}-G_{ij}^{(k)})\|_{L_{r}}=\|\phi G_{ik}\sum_{l}^{(k)}H_{kl}G_{lj}^{(k)}\|_{L_{r}}$
and applying (6.1). We omit the details.
To prove (6.11), we start from the decomposition (6.7), which gives
$Y_{i}G_{ii}=G_{ii}H_{ii}+\frac{1}{n}\sum_{k}G_{ki}G_{ik}-G_{ii}\sum_{k\neq l}^{(i)}H_{ik}G_{kl}^{(i)}H_{li}-G_{ii}\sum_{k}^{(i)}(H_{ik}^{2}-n^{-1})G_{kk}^{(i)}.$ (6.15)
We proceed by bounding each term in this expression separately. From the definition of $H_{ii}$, we first have
$\|\phi G_{ii}H_{ii}\|_{L_{r}}\leq\frac{2}{\sqrt{np}}\|\widetilde{H}_{ii}\|_{L_{r}}\leq\frac{2Kr^{\frac{1}{\alpha}}}{\sqrt{np}}.$
Next, by Ward identity (6.8), we obtain
$\phi\frac{1}{n}\left|\sum_{k}G_{ki}G_{ik}\right|\leq\phi\frac{1}{n}\sum_{k}|G_{ik}|^{2}=\phi\frac{\operatorname{Im}G_{ii}}{n\eta}\leq\frac{2}{n\eta}.$
For the third term on the right-hand side of (6.15), we use the nesting property of conditional $L_{r}$ norms to derive
$\Big{\|}\phi G_{ii}\sum_{k\neq l}^{(i)}H_{ik}G_{kl}^{(i)}H_{li}\Big{\|}_{L_{r}}$ $\leq 2\left\|\phi^{(i)}\Big{\|}\sum_{k\neq l}^{(i)}H_{ik}G_{kl}^{(i)}H_{li}\Big{\|}_{L_{r}|H^{(i)}}\right\|_{L_{r}}$ (6.16)
$=\frac{2}{np}\left\|\phi^{(i)}\Big{\|}\sum_{k\neq l}^{(i)}\widetilde{H}_{ik}G_{kl}^{(i)}\widetilde{H}_{li}\Big{\|}_{L_{r}|H^{(i)}}\right\|_{L_{r}}.$ (6.17)
Applying (6.3), we further obtain
$\phi^{(i)}\Big{\|}\sum_{k\neq l}^{(i)}\widetilde{H}_{ik}G_{kl}^{(i)}\widetilde{H}_{li}\Big{\|}_{L_{r}|H^{(i)}}$ (6.18)
$\qquad\leq\phi^{(i)}C_{\alpha}K^{2}\max\left\{r^{\max\{\frac{2}{\alpha},2\}}\|G_{0}^{(i)}\|_{\max},rp\|G_{0}^{(i)}\|_{1,\infty},\sqrt{r}p\|G_{0}^{(i)}\|_{F}\right\},$ (6.19)
where the matrix $G_{0}^{(i)}$ is obtained from $G^{(i)}$ by setting all diagonal entries to zero. Observe that $\phi^{(i)}\|G_{0}^{(i)}\|_{\max}\leq 2$ and by Cauchy-Schwartz inequality and (6.8), we have
$\phi^{(i)}\|G_{0}^{(i)}\|_{1,\infty}$ $=\phi^{(i)}\max_{k\neq i}\sum_{j}^{(i)}|G_{kj}^{(i)}|$ (6.20)
$\leq\phi^{(i)}\sqrt{n}\max_{k\neq i}\sqrt{\sum_{j}^{(i)}|G_{kj}^{(i)}|^{2}}=\phi^{(i)}\sqrt{n}\max_{k\neq i}\sqrt{\frac{\operatorname{Im}G_{kk}^{(i)}}{\eta}}\leq 2\frac{\sqrt{n}}{\sqrt{\eta}}.$ (6.21)
It follow from (6.8) that
$\phi^{(i)}\|G_{0}^{(i)}\|_{F}=\phi^{(i)}\sqrt{\sum_{k\neq j}^{(i)}|G_{kj}^{(i)}|^{2}}\leq\phi^{(i)}\sqrt{\sum_{k}^{(i)}\frac{\operatorname{Im}G_{kk}^{(i)}}{\eta}}\leq 2\frac{\sqrt{n}}{\sqrt{\eta}}.$ (6.22)
Hence, we arrive at
$\Big{\|}\phi G_{ii}\sum_{k\neq l}^{(i)}H_{ik}G_{kl}^{(i)}H_{li}\Big{\|}_{L_{r}}\leq 4C_{\alpha}K^{2}\max\left\{\frac{r^{\max\{\frac{2}{\alpha},2\}}}{np},\frac{r}{\sqrt{n\eta}}\right\}.$ (6.23)
For the third term on the right-hand side of (6.15), similar calculation using (6.2) yields
$\Big{\|}\phi G_{ii}\sum_{k}^{(i)}(H_{ik}^{2}-n^{-1})G_{kk}^{(i)}\Big{\|}_{L_{r}}$ $\leq\frac{2}{np}\left\|\Big{\|}\phi^{(i)}\sum_{k}^{(i)}(\widetilde{H}_{ik}^{2}-p)G_{kk}^{(i)}\Big{\|}_{L_{r}|H^{(i)}}\right\|_{L_{r}}$ (6.24)
$\leq C_{\alpha}K^{2}\frac{4}{np}\max\left\{r^{\max\{\frac{2}{\alpha},1\}},\sqrt{r}\sqrt{np}\right\}$ (6.25)
$=4C_{\alpha}K^{2}\max\left\{\frac{r^{\max\{\frac{2}{\alpha},1\}}}{np},\frac{\sqrt{r}}{\sqrt{np}}\right\}.$ (6.26)
Combining the above estimates, we obtain
$\max_{i}\|\phi Y_{i}G_{ii}\|_{L_{r}}\leq$ $\frac{2Kr^{\frac{1}{\alpha}}}{\sqrt{np}}+\frac{2}{n\eta}+4C_{\alpha}K^{2}\max\left\{\frac{r^{\max\{\frac{2}{\alpha},2\}}}{np},\frac{r}{\sqrt{n\eta}}\right\}$ (6.27)
$+4C_{\alpha}K^{2}\max\left\{\frac{r^{\max\{\frac{2}{\alpha},1\}}}{np},\frac{\sqrt{r}}{\sqrt{np}}\right\}.$ (6.28)
By our supposition $r^{\max\{\frac{1}{\alpha},1\}}\leq\sqrt{np}$, we bound
$\frac{r^{\max\{\frac{2}{\alpha},2\}}}{np}\leq\frac{r^{\max\{\frac{1}{\alpha},1\}}}{\sqrt{np}}.$
Since $r\geq 2$, $C_{\alpha}\geq 1$, $K\geq 1$ and $n\eta\geq 1$, we obtain a simplified bound
$\max_{i}\|\phi Y_{i}G_{ii}\|_{L_{r}}\leq 12C_{\alpha}K^{2}\max\left\{\frac{r^{\max\{\frac{1}{\alpha},1\}}}{\sqrt{np}},\frac{r}{\sqrt{n\eta}}\right\}$ (6.29)
and this completes the proof of (6.11). ∎
Now we proceed with the proof of Theorem 12. Since $G_{ij}$, $s$, $m$ and $\zeta$ are $n^{2}$-Lipschitz on $\mathbf{S}$, it suffices to prove the following result for the spectral parameter $z=E+\mathrm{i}\eta$ (with fixed $E$) being in the lattice set $\{z_{0},z_{1},\ldots,z_{\mathcal{K}}\}$, where $z_{k}:=E+\mathrm{i}\eta_{k}$ with $\eta_{k}:=1-kn^{-3}$. Here $\mathcal{K}:=\max\{k\in\mathbb{N}:1-kn^{-3}\geq n^{-1}\}$ and $\mathcal{K}\leq n^{3}-n^{2}$. Recall $\zeta=\zeta(n,r,p,\eta,\alpha)$ from (6.4). We denote $\zeta(n,r,p,\eta_{k},\alpha):=\zeta_{k}$ below.
PROPOSITION A. 1.
Assume $n\geq 10$, $r\geq 2$ and $r^{\max\{\frac{1}{\alpha},1\}}\leq\sqrt{np}$. For any $t\geq 1$ satisfying $20\zeta_{k}t\leq 1$ for every $k=0,1,\ldots,\mathcal{K}$, we have
$\mathbb{P}\left(\max_{i,j}|G_{ij}(z_{k})-m(z_{k})\delta_{ij}|>10\zeta_{k}t\right)\leq 3kn^{3}t^{-2r}.$ (6.30)
The proof of Proposition 1 follows the standard stochastic continuity argument as seen in [10, 24]. For completeness, we include the proof in Appendix D.
The proof of Theorem 12 follows from Proposition 1, the Lipschitz continuity of $G_{ij}$ and $m$, $\zeta$ is a decreasing function of $\eta$, and $\frac{1}{(n\eta)^{2}}\leq\zeta t$, and a union bound on the lattice set, following the same arguments as in [24]. We omit the details.
6.2 Proof of Theorem 11
Proof of Theorem 11.
Note that $\|BX\|^{2}=X^{T}B^{T}BX$. To apply Corollary 2 to matrix $A=B^{T}B$ and random vector $X$, note that $(B^{T}B)_{ij}=\langle B_{i},B_{j}\rangle$ and the expectation can be computed as
$\mathbb{E}\|BX\|^{2}=\mathbb{E}X^{T}B^{T}BX=\sum_{i=1}^{n}(B^{T}B)_{ii}\mathbb{E}X_{i}^{2}=p\operatorname{tr}(B^{T}B)=p\|B\|_{F}^{2}.$
Moreover, the relevant matrix norms can be expressed as
$p^{2}\sum_{i\neq j}(B^{T}B)_{ij}^{2}+p\sum_{i=1}^{n}(B^{T}B)_{ii}^{2}$ $=p^{2}\|B^{T}B\|_{F}^{2}+p\sum_{i=1}^{n}\|B_{i}\|^{4},$
$\gamma_{1,\infty}(B^{T}B)$ $=\max_{1\leq i\leq n}\Big{\{}p\sum_{j\neq i}|\langle B_{i},B_{j}\rangle|,\|B_{i}\|^{2}\Big{\}},$
$\|B^{T}B\|_{\max}$ $=\max_{i,j}|\langle B_{i},B_{j}\rangle|=\|B^{T}\|_{2,\infty}^{2}.$
Hence, applying the tail bound in Corollary 2 with
$t=$ $\;\sqrt{s}\cdot K^{2}\sqrt{p^{2}\|B^{T}B\|_{F}^{2}+p\sum_{i}\|B_{i}\|^{4}}+s\cdot K^{2}\max_{1\leq i\leq n}\Big{\{}p\sum_{j\neq i}|\langle B_{i},B_{j}\rangle|,\|B_{i}\|^{2}\Big{\}}$
$+s^{\max\{\frac{2}{\alpha},2\}}\cdot K^{2}\|B^{T}\|_{2,\infty}^{2},$
we have
$\mathbb{P}\left(|\|BX\|^{2}-{p}\|B\|_{F}^{2}|>t\right)\leq C^{\prime}\exp(-cs).$
Acknowledgments
We thank Fanny Augeri, Antti Knowles, Stanislav Minsker, Roman Vershynin, and Shuheng Zhou for helpful discussions. K. Wang is partially supported by Hong Kong RGC grant GRF 16304222 and ECS 26304920. Y. Zhu was partially supported by NSF-Simons Research Collaborations on the Mathematical and Scientific Foundations of Deep Learning and an AMS-Simons Travel Grant. This material is based upon work supported by the Swedish Research Council under grant no. 2021-06594 while Y. Zhu was in residence at Institut Mittag-Leffler in Djursholm, Sweden during the Fall of 2024.
Appendix A Auxiliary inequalities
LEMMA A.1 (Gibbs’ inequality [32]).
Suppose that $(p_{1},\dots,p_{n}),(q_{1},\dots,q_{n})$ are discrete probability measures, i.e., $0\leq p_{i},q_{i}\leq 1$, $\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}q_{i}=1$. Then
$-\sum_{i=1}^{n}p_{i}\log(p_{i})\leq-\sum_{i=1}^{n}p_{i}\log q_{i}.$ (A.1)
LEMMA A.2.
For positive $1\leq a\leq b$, there is
$a^{a}b^{b}<(a-1)^{(a-1)}(b+1)^{(b+1)}.$
In general, if we have positive integers $L$ and $x_{1},x_{2},\dots,x_{k}\geq L\geq 1$ with $x_{1}+\dots+x_{k}=S\geq kL$, then the maximal of $\prod_{i=1}^{k}x_{i}^{x_{i}}$ is taken when all but one $x_{i}$’s are $L$, i.e.,
$\prod_{i=1}^{k}x_{i}^{x_{i}}\leq L^{(k-1)L}(S-(k-1)L)^{(S-(k-1)L)}.$
Proof.
We first prove the first inequality, which is equivalent to show that
$\frac{a^{a}}{(a-1)^{a-1}}<\frac{(b+1)^{b+1}}{b^{b}}.$
Note that the function $f(x)=(x+1)(1+\frac{1}{x})^{x}$ is strictly in creasing on $(0,+\infty)$. We know the inequality holds as $f(a-1)<f(b)$.
In general, suppose $x_{1},\dots,x_{k}$ maximize $\prod_{i=1}^{k}x_{i}^{x_{i}}$ and there are at least two $x_{i},x_{j},i\neq j$ such that $L<x_{i}\leq x_{j}$. We replace $x_{i},x_{j}$ with $x_{i}-1,x_{j}+1$. Then the constraints still hold, and by the first part of the lemma we obtain a larger value of $\prod_{i=1}^{k}x_{i}^{x_{i}}$, which leads to a contradiction. Therefore, the maximum is taken where all but one $x_{i}$’s are $L$. ∎
Appendix B Generalization to noncentered random variables
In this section, we briefly discuss the extension of our results to the concentration of quadratic forms of noncentered random variables. Let $Y_{1},\cdots,Y_{n}$ be independent random variables, and consider a symmetric, diagonal-free matrix $A$ for simplicity.
Denote $X_{i}=Y_{i}-\mathbb{E}(Y_{i})$. We rewrite
$Y^{T}AY=\sum_{i\neq j}a_{ij}Y_{i}Y_{j}=\sum_{i\neq j}a_{ij}(X_{i}+\mathbb{E}(Y_{i}))(X_{j}+\mathbb{E}(Y_{j})).$ (B.1)
Through straightforward calculation, we obtain
$Y^{T}AY-\mathbb{E}(Y^{T}AY)=X^{T}AX+2\sum_{i}\left(\sum_{j}a_{ij}\mathbb{E}(Y_{j})\right)X_{i}:=X^{T}AX+2\sum_{i=1}^{n}b_{i}X_{i}.$
Note that $\mathrm{Var}(Y^{T}AY)=\mathrm{Var}(X^{T}AX)+\mathrm{Var}(2\sum_{i}b_{i}X_{i})$ since $X^{T}AX$ and $\sum_{i=1}^{n}b_{i}X_{i}$ are uncorrelated.
The concentration inequality for $Y^{T}AY-\mathbb{E}(Y^{T}AY)$ can be obtained by considering those of $X^{T}AX$ and $\sum_{i}b_{i}X_{i}$ separately. Assume the moments of $Y_{i}$ satisfy $\mathbb{E}|Y_{i}|^{r}\leq p_{i}(Kr)^{r}$ for $r\geq 1$. Then from the inequality $\|X_{i}\|_{L_{r}}=\|Y_{i}-\mathbb{E}(Y_{i})\|_{L_{r}}\leq 2\|Y_{i}\|_{L_{r}}$, we deduce that the $X_{i}$’s are independent centered random variables satisfying $\mathbb{E}|X_{i}|^{r}\leq p_{i}(2Kr)^{r}$ for $r\geq 2$. Consequently, the conclusion of Theorem 1 holds for $X^{T}AX$. Moreover, $\sum_{i=1}^{n}b_{i}X_{i}$ is a sum of independent random variables. Applying Corollary 1 with $\alpha=2$, we find that the tail bound for $\sum_{i=1}^{n}b_{i}X_{i}$ is a mixture of sub-gaussian and sub-exponential tails. For the sub-exponential tail, we can use the estimate that
$\max_{i}|b_{i}|=\max_{i}|\sum_{j}a_{ij}\mathbb{E}(Y_{j})|\leq K\max_{i}\sum_{j}|a_{ij}|p_{j}=K\gamma_{1,\infty}.$
Based on the preceding discussion, we now present a generalization of Theorem 1 for $\sum_{i\neq j}a_{ij}Y_{i}Y_{j}$, where the random variables $Y_{i}$ are not necessarily centered.
THEOREM 13.
Let $Y=(Y_{1},\cdots,Y_{n})^{T}$ be a random vector with independent sparse sub-exponential entries, i.e., each $Y_{i}$ satisfies $\mathbb{E}|Y_{i}|^{r}\leq p_{i}(Kr)^{r}$ for all $r\geq 1$, where $p_{i}\in(0,1)$ and $K>0$. Let $A$ be an $n\times n$ diagonal-free matrix. Then there exists an absolute constant $C>0$ such that for any $t>0$,
$\mathbb{P}\left(\left|Y^{T}AY-\mathbb{E}(Y^{T}AY)\right|\geq t\right)$ (B.2)
$\leq C^{\prime}\exp\left(-C\min\left\{\frac{t^{2}}{K^{4}(\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}+\sum_{i}b_{i}^{2}p_{i})},\frac{t}{K^{2}\gamma_{1,\infty}},\left(\frac{t}{K^{2}\|A\|_{\max}}\right)^{1/2}\right\}\right),$ (B.3)
where $b_{i}:=\max\{\sum_{j\neq i}a_{ij}\mathbb{E}(Y_{j}),\sum_{j\neq i}a_{ji}\mathbb{E}(Y_{j})\}$ and $\gamma_{1,\infty}$ is defined in (1.3).
Appendix C Proof of Theorem 10
We begin the proof of Theorem 10 by showing that, with high probability, the eigenvalues of $H$ lie within the spectral domain $\mathbf{S}$. This is done by the spectral norm concentration inequality from [30] with a truncation argument.
LEMMA C.1.
There is an absolute constant $C>0$ depending such that for any $D>0$, if $p\geq\frac{K^{2}e^{2(D+1)}(\log n)^{1+\frac{2}{\alpha}}}{n}$, with probability at least $1-2n^{-D}$,
$\|\widetilde{H}\|\leq C(1+\sqrt{D})\sqrt{np}.$ (C.1)
Proof.
Since the entries of $\widetilde{H}$ satisfy $\mathbb{E}|\widetilde{H}_{ij}|^{r}\leq K^{r}r^{\frac{r}{\alpha}}$ for all $r\geq 1$, applying Markov’s inequality with a union bound yields
$\mathbb{P}\left(|\widetilde{H}_{ij}|>2e^{D+1}K(\log n)^{\frac{1}{\alpha}}\right)\leq n^{-2-D},$ (C.2)
$\mathbb{P}\left(\|\widetilde{H}\|_{\infty}>2e^{D+1}K(\log n)^{\frac{1}{\alpha}}\right)\leq n^{-D}.$ (C.3)
Denote $B:=2e^{D+1}K(\log n)^{\frac{1}{\alpha}}$. Define $\check{H}_{ij}=\widetilde{H}_{ij}\mathbf{1}\{|\widetilde{H}_{ij}|\leq B\}$. Then $|\check{H}_{ij}-\mathbb{E}\check{H}_{ij}|\leq 2B$. [30, Theorem 4.8] implies
$\|\check{H}-\mathbb{E}\check{H}\|\leq C\left(\sqrt{np}+B\sqrt{\log n}+t\right)$ (C.4)
with probability at least $1-\exp(-\frac{t^{2}}{CB^{2}})$. Taking $t=\sqrt{CD\log n}B$ implies with probability at least $1-n^{-D}$, $\|\check{H}-\mathbb{E}\check{H}\|\leq C(1+\sqrt{D})(\sqrt{np}+B\sqrt{\log n})$. On the other hand, since $\mathbb{E}\check{H}$ is of rank-1,
$\|\mathbb{E}\check{H}\|$ $=n|\mathbb{E}\check{H}_{ij}|=n|\mathbb{E}\check{H}_{ij}-\mathbb{E}\widetilde{H}_{ij}|=n|\mathbb{E}\widetilde{H}_{ij}\mathbf{1}\{|\widetilde{H}_{ij}|>B\}|$ (C.5)
$\leq n\sqrt{\mathbb{E}\widetilde{H}_{ij}^{2}}\sqrt{\mathbb{P}(|\widetilde{H}_{ij}|>B)}\leq Cn\sqrt{p}n^{-1-D/2}\leq C\sqrt{np}.$ (C.6)
Therefore with probability at least $1-2n^{-D}$,
$\|\widetilde{H}\|=\|\check{H}\|$ $\leq\|\check{H}-\mathbb{E}\check{H}\|+\|\mathbb{E}\check{H}\|$ (C.7)
$\leq C(1+\sqrt{D})(\sqrt{np}+Ke^{D+1}(\log n)^{\frac{1}{2}+\frac{1}{\alpha}})$ (C.8)
$\leq 2C(1+\sqrt{D})\sqrt{np},$ (C.9)
where in the last inequality, we use the assumption that $p\geq\frac{K^{2}e^{2(D+1)}(\log n)^{1+2/\alpha}}{n}$. ∎
We work on the conclusion of Lemma C.1. This result implies that for each eigenvalue $\lambda_{i}$ of $H$, $z_{i}:=\lambda_{i}+\mathrm{i}\eta$, where $\eta=LC_{\alpha}^{2}K^{4}\frac{\log^{2}n}{n}$, lies in a spectral domain
$\hat{\mathbf{S}}\equiv\hat{\mathbf{S}}(D):=\{E+\mathrm{i}\eta\in\mathbb{C}:|E|\leq C(1+\sqrt{D}),n^{-1}<\eta\leq 1\}.$
The domain $\hat{\mathbf{S}}$ is a slight modification of the spectral domain $\mathbf{S}$ defined in (3.2). An inspection of the proof of Theorem 8 reveals that there exist constants $L\equiv L(\delta,D)$ and $n_{0}\equiv n_{0}(\delta,D)$ such that for $p\geq LC_{\alpha}^{2}K^{4}\frac{(\log n)^{\max\{2,\frac{2}{\alpha}\}}}{n}$, we have
$\mathbb{P}\left(\max_{i,j}\left|G_{ij}(z)-m(z)\delta_{ij}\right|\leq\delta\right)\geq 1-n^{-D}.$ (C.10)
for any $z=E+\mathrm{i}\eta\in\hat{\mathbf{S}}$ with $\eta\geq LC_{\alpha}^{2}K^{4}\frac{\log^{2}n}{n}$, provided $n\geq n_{0}$. The only modification needed in the proof occurs in the final step, where we take a union bound over the lattice set $\hat{\mathbf{S}}\cap n^{-3}{\mathbb{Z}}^{2}$. The size of this lattice set depends on $D$, and we omit the detailed analysis.
Therefore, taking $\delta=1/2$ in (C.10) and using the bound $|m|\leq 1$, we have
$\frac{3}{2}\geq\operatorname{Im}G_{kk}(z_{i})=\sum_{j=1}^{n}\frac{\eta}{(\lambda_{j}-\lambda_{i})^{2}+\eta^{2}}|u_{j}(k)|^{2}\geq\frac{1}{\eta}|u_{i}(k)|^{2},$ (C.11)
which yields $|u_{i}(k)|\leq\sqrt{\frac{2}{\eta}}=\sqrt{2L}C_{\alpha}K^{2}\frac{\log n}{\sqrt{n}}$. This completes the proof.
Appendix D Proof of Proposition 1
We now prove Proposition 1, following closely the proof of Proposition 4.5 from [24]. We outline the main steps below. First, we recall the stability estimate from Lemma 4.4 of [24]: For $z\in\mathbf{S}$, let $m,\tilde{m}$ be the solutions of the equation $x^{2}+zx+1=0$. If $s$ satisfies $s^{2}+zs+1=r$, then
$\min\{|s-m|,|s-\tilde{m}|\}\leq\sqrt{|r|}.$ (D.1)
From Lemma 6.1 and the definition of $\zeta_{k}$, we have
$\max\left\{\max_{i\neq j}\|\phi G_{ij}(z_{k})\|_{L_{r}},\max_{i,j\neq k}\|\phi(G_{ij}-G_{ij}^{(k)})(z_{k})\|_{L_{r}},\max_{i}\|\phi Y_{i}G_{ii}(z_{k})\|_{L_{r}}\right\}\leq\zeta_{k}^{2}.$ (D.2)
Define the events
$\Omega_{k}:=\{|s(z_{k})-m(z_{k})|\leq\zeta_{k}t\}\quad\text{and}\quad\Xi_{k}:=\{\Gamma(z_{k})\leq 3/2\}$
for $k=0,1,\ldots,\mathcal{K}$. We proceed with an induction on $k$.
For $k=0$, we have $\eta_{0}=1$ and $\Gamma(z_{0})\leq 1$ using the trivial bound $\|G(z_{0})\|\leq 1$. Hence, $\phi(z_{0})=1$ and $\mathbb{P}(\Xi_{0})=1$. From the $1+zs+s^{2}=\frac{1}{n}\sum_{i}G_{ii}Y_{i}$ and (D.2), we obtain
$\|1+z_{0}s(z_{0})+s(z_{0})^{2}\|_{L_{r}}\leq\max_{i}\|\phi Y_{i}G_{ii}(z_{0})\|_{L_{r}}\leq\zeta_{0}^{2}.$
By Markov’s inequality, we further get
$\mathbb{P}\left(|1+z_{0}s(z_{0})+s(z_{0})^{2}|>\zeta_{0}^{2}t^{2}\right)\leq t^{-2r}.$
It follows from (D.1) that
$\mathbb{P}\left(\min\{|s(z_{0})-m(z_{0})|,|s(z_{0})-\tilde{m}(z_{0})|\}>\zeta_{0}t\right)\leq t^{-2r}.$ (D.3)
From $\operatorname{Im}\tilde{m}(z_{0})<-1$, $\operatorname{Im}s>0$ and $\zeta_{0}t<1/10$, we conclude that
$\mathbb{P}(\Omega_{0}^{c})=\mathbb{P}\left(|s(z_{0})-m(z_{0})|>\zeta_{0}t\right)\leq t^{-2r}.$ (D.4)
Likewise, (D.2), together with Markov’s inequality, yields
$\max_{i}\mathbb{P}(|Y_{i}(z_{0})G_{ii}(z_{0})|>\zeta_{0}t)\leq t^{-2r}\quad\text{and}\quad\max_{i\neq j}\mathbb{P}(|G_{ij}(z_{0})|>\zeta_{0}t)\leq t^{-2r}.$ (D.5)
These estimates, combined with $|G_{ii}-m|\leq|(s-m)G_{ii}|+|Y_{i}G_{ii}|$ (see (4.29) from [24]), lead to
$\max_{i}\mathbb{P}(|G_{ii}(z_{0})-m(z_{0})|>2\zeta_{0}t)\leq 2t^{-2r}.$ (D.6)
Hence, by a union bound, we get
$\mathbb{P}\left(\max_{i,j}|G_{ij}(z_{0})-m(z_{0})\delta_{ij}|>2\zeta_{0}t\right)\leq(n^{2}+2n)t^{-2r}<3n^{2}t^{-2r}.$ (D.7)
For $k\geq 1$, we show that conditioning on $\Omega_{k-1}\cap\Xi_{k-1}$, the event $\Omega_{k}\cap\Xi_{k}$ holds with high probability. Define a threshold index $\widetilde{\mathcal{K}}:=\min\{k\in\mathbb{N}:k\leq\mathcal{K},|m(z_{k})-\tilde{m}(z_{k})|\leq 5\zeta_{k}t\}.$ For $k<\widetilde{\mathcal{K}}$, $|m(z_{k})-\tilde{m}(z_{k})|>5\zeta_{k}t$, and we must sure that $s(z_{k})$ is close to $m(z_{k})$ rather than $\tilde{m}(z_{k})$. In contrast, for $k\geq\widetilde{\mathcal{K}}$, the distinction between $m(z_{k})$ and $\tilde{m}(z_{k})$ becomes insignificant, leading to a simpler argument. The majority of the proof follows steps analogous to those used for the base case $k=0$. We provide a sketch of the key arguments below.
Case I. For $1\leq k<\widetilde{\mathcal{K}}$, since $\Gamma(z)$ is $n^{2}$-Lipschitz, we have $\phi(z_{k})=1$ on $\Xi_{k-1}$. Following the derivation of (D.3), we obtain
$\mathbb{P}\left(\phi(z_{k})\min\{|s(z_{k})-m(z_{k})|,|s(z_{k})-\tilde{m}(z_{k})|\}>\zeta_{k}t\right)\leq t^{-2r}.$
Since $\zeta$ is a decreasing function of $\eta$ and $m,s$ are $n^{2}$-Lipschitz, on $\Omega_{k-1}$ we have
$|s(z_{k})-m(z_{k})|\leq|s(z_{k-1})-m(z_{k-1})|+\frac{2n^{-3}}{\eta^{2}}\leq\zeta_{k-1}+\frac{2}{n^{3}\eta}<2\zeta_{k}t.$
This implies that $\min\{|s(z_{k})-m(z_{k})|,|s(z_{k})-\tilde{m}(z_{k})|=|s(z_{k})-m(z_{k})|$. Therefore,
$\mathbb{P}(\Omega_{k-1}\cap\Xi_{k-1}\cap\Xi_{k}^{2})\leq t^{-2r}.$
Similarly to (D.5) and (D.6), the following estimates hold:
$\max_{i}\mathbb{P}(\Omega_{k-1}\cap\Xi_{k-1}\cap\{|G_{ii}(z_{k})-m(z_{k})|>2\zeta_{k}t\})\leq 2t^{-2r},$ (D.8)
$\max_{i\neq j}\mathbb{P}(\Xi_{k-1}\cap\{|G_{ij}(z_{k})|>\zeta_{k}t\})\leq t^{-2r},$ (D.9)
$\max_{i,j\neq l}\mathbb{P}(\Xi_{k-1}\cap\{|G_{ij}(z_{k})-G_{ij}^{(l)}(z_{k})|>\zeta_{k}t\})\leq t^{-2r}.$ (D.10)
A union bound, together with the estimate $|m|\leq 1$, implies that
$\mathbb{P}(\Omega_{k-1}\cap\Xi_{k-1}\cap\{\Gamma(z_{k})>1+2\zeta_{k}t\})\leq(2n+n^{2}+n^{3})t^{-2r}\leq 2n^{3}t^{-2r}.$
Noting that $\zeta_{k}t\leq 1/20$, we obtain
$\mathbb{P}(\Omega_{k-1}\cap\Xi_{k-1}\cap\Xi_{k}^{c})\leq 2n^{3}t^{-2r}$
and consequently,
$\mathbb{P}\left(\Omega_{k-1}\cap\Xi_{k-1}\cap(\Omega_{k}\cap\Xi_{k})^{c}\right)\leq(2n^{3}+1)t^{-2r}.$
Since $\mathbb{P}((\Omega_{0}\cap\Xi_{0})^{c})\leq t^{-2r}$, a standard induction argument (see, for instance, (5.17) from [10]) shows that
$\mathbb{P}((\Omega_{k}\cap\Xi_{k})^{c})\leq t^{-2r}+k(2n^{3}+1)t^{-2r}.$
Finally, similar to (D.7), we arrive at
$\mathbb{P}\left(\max_{i,j}|G_{ij}(z_{k})-m(z_{k})\delta_{ij}|>2\zeta_{k}t\right)\leq t^{-2r}+k(2n^{3}+1)t^{-2r}+(2n+n^{2})t^{-2r}\leq 3kn^{3}t^{-2r}.$
Case II. For $\widetilde{\mathcal{K}}\leq k\leq\mathcal{K}$, given that $|\tilde{m}(z_{k})-m(z_{k})|<5\zeta_{k}t$, we have the following estimates:
$\mathbb{P}(\Xi_{k-1}\cap\{|s(z_{k})-m(z_{k})|>6\zeta_{k}t)\leq t^{-2r},$ (D.11)
$\max_{i}\mathbb{P}(\Xi_{k-1}\cap\{|G_{ii}(z_{k})-m(z_{k})|>10\zeta_{k}t\})\leq 2t^{-2r},$ (D.12)
$\max_{i\neq j}\mathbb{P}(\Xi_{k-1}\cap\{|G_{ij}(z_{k})|>\zeta_{k}t\})\leq t^{-2r}.$ (D.13)
Since $|m|\leq 1$ and $20\zeta_{k}t\leq 1$, applying a union bound yields
$\mathbb{P}(\Xi_{k-1}\cap\Xi_{k}^{c})\leq(2n+n^{2})t^{-2r}.$
Using the bound $\mathbb{P}(\Xi_{\widetilde{K}-1}^{c})\leq 3(\widetilde{K}-1)n^{3}t^{-2r}$ from Case I and an induction argument, we obtain
$\mathbb{P}(\Xi_{k}^{c})\leq 3(\widetilde{K}-1)n^{3}t^{-2r}+(k-\widetilde{K})(2n+n^{3})t^{-2r}.$
Finally, following an argument similar to (D.7), we conclude
$\mathbb{P}\left(\max_{i,j}|G_{ij}(z_{k})-m(z_{k})\delta_{ij}|>10\zeta_{k}t\right)\leq 3kn^{3}t^{-2r}.$